### v4

# A radial scalar product for Kerr quasinormal modes

Abstract
A scalar product for quasinormal mode solutions to Teukolsky’s homogeneous radial equation is presented. Evaluation of this scalar product can be performed either by direct integration, or by evaluation of a confluent hypergeometric functions. The related scalar product will be useful for better understanding analytic solutions to Teukolsky’s radial equation, particularly the quasi-normal modes, their potential spatial completeness, and whether the quasi-normal mode overtone excitations may be estimated by spectral decomposition rather than fitting. With that motivation, the scalar product is applied to confluent Heun polynomials where it is used to derive their peculiar orthogonality and eigenvalue properties. A potentially new relationship is derived between the confluent Heun polynomials’ scalar products and eigenvalues. Using these results, it is shown for the first time that Teukolsky’s radial equation (and perhaps similar confluent Heun equations) are, in principle, exactly tri-diagonalizable. To this end, “canonical” confluent Heun polynomials are conjectured.
I Introduction
The aftermath of binary black hole (BBH) merger is thought to be a spinning and perturbed black hole (BH) that rings down like a struck, and shimmering gong Vishveshwara (1970); Abbott et al. (2016a, b, 2021a, 2021b, 2019a). The corresponding gravitational radiation encodes information about the progenitor BBH system (i.e. masses, spins, momenta and environment), and about the true nature of gravity Kamaretsos et al. (2012); London et al. (2014); Hughes et al. (2019); Berti et al. (2006a, 2009); Maselli et al. (2017); Kokkotas and Schmidt (1999); Evstafyeva et al. (2023); Arun et al. (2022); Pacilio and Bhagwat (2023). Ongoing and planned experiments endeavor to detect and learn from that radiation with increasing precision Bailes et al. (2021); Sathyaprakash et al. (2012); Berti et al. (2006a); Vallisneri et al. (2015); Karnesis et al. (2022); Arun et al. (2022). At all stages, those efforts are guided by our mathematical understanding of BHs, their gravitational waves, and the special functions best suited to them Abbott et al. (2017, 2016c); Misner et al. (1973); Maggiore (2007); Newman and Penrose (1966, 1962); Teukolsky (1972); Whiting (1989). Recently, it has become clear that there are limitations in our mathematical understanding of ringdown radiation, and that these limitations may actively confound the interpretation of numerical and experimental data Giesler et al. (2019); McWilliams (2019); Carullo et al. (2019); Ota and Chirenti (2020); Bhagwat et al. (2020); Forteza et al. (2020); Leong et al. (2023); Khera et al. (2023); Franchini and Völkel (2023); Nee et al. (2023); Cotesta et al. (2022); Baibhav et al. (2023).
The purpose of this article is to introduce a framework via a radial scalar-product that may offer a deeper, practical understanding of ringdown “overtones,” and potentially of gravitational waves from compact binaries more broadly. The primary technical focus of this work is the vector space structure of the eigensolutions to Teukolskys equation for Quasi-Normal Modes (QNMs), considered after the imposition of physical boundary conditions.
To date, the LIGO-Virgo-Kagra collaboration (LVK) has confidently observed gravitational radiation from $83$ BBH mergers; among them, the ringdown radiation of $31$ have been compared against gravitational wave signal models (theory predictions) to test for physics beyond General Relativity (GR) Abbott et al. (2021a, b, 2019a, c, 2019b, d, e). Thus far, all BBH remnants have been found to be consistent with spinning (i.e. Kerr) BHs Abbott et al. (2019b, 2021d, 2021e).
Planned gravitational wave detectors such as the Einstein Telescope (ET), Cosmic Explorer (CE) and the Laser Interferometer Space Antenna (LISA) are, for example, expected to detect signal to noise ratios of $\sim 10^{3}$ (ET & CE) and $\sim 10^{5}$ (LISA) from GW150914-like [^1] This example refers to BBHs that have similar configurations to that of the first gravitational wave detection: approximately non-spinning progenitors and having a luminosity distance of $\sim 410~\mathrm{Mpc}$ Abbott et al. (2016c, a). BBHs of $10^{2}~\mathrm{M}_{\odot}$ and $10^{6}~\mathrm{M}_{\odot}$, respectively Berti et al. (2016); Mishra et al. (2010); Robson et al. (2019). Such detections are expected to dramatically increase the potential for ringdown data analyses to verify proposed environmental, dark matter and quantum effects Klein et al. (2016); Franchini and Völkel (2023); Macedo et al. (2013); Berti et al. (2006a); Barausse et al. (2015); Zhao et al. (2023). However, the correct interpretation of analyses involving ringdown is contingent upon our understanding of ringdown’s underlying mathematical structure Berti et al. (2003); Berti and Klein (2014); Berti et al. (2006b); London (2020); London et al. (2014); London (2023); London and Fauchon-Jones (2019); Husa et al. (2016); Hamilton et al. (2021, 2023); Cotesta et al. (2020); Pompili et al. (2023); Ossokine et al. (2020); Hughes et al. (2019); Lagos and Hui (2022).
Here, ringdown will refer to linear combinations of BH QNMs Press (1971); Leaver (1985); Press and Teukolsky (1973). While the gravitational waves from BBH postmerger will refer to all linear and nonlinear contributions London et al. (2014), the mathematical structure of ringdown is that of the QNMs. Due to their astrophysical relevance, the QNMs of Kerr BHs are perhaps the most studied Berti et al. (2003); Berti and Klein (2014); Berti et al. (2006b); London (2020); London et al. (2014); London (2023); London and Fauchon-Jones (2019); Husa et al. (2016); Hamilton et al. (2021, 2023); Cotesta et al. (2020); Pompili et al. (2023); Ossokine et al. (2020); Hughes et al. (2019); Ghosh et al. (2018); Yang et al. (2012, 2013); Zimmerman and Mark (2016); Lagos and Hui (2022).
The Kerr QNMs emerge within the study of Einstein’s equations linearized around the Kerr metric Vishveshwara (1970); Press and Teukolsky (1973); Leaver (1985). Within that topic, a central result is that each single QNM solution is a function of time (a damped sinusoid), times a function of spherical polar angles (a spheroidal harmonic), and times a function of radial distance (a confluent Heun function) Press and Teukolsky (1973); Dariescu et al. (2021); Fiziev (2010); Hatsuda (2021).
The time domain’s sinusoidal damping is determined by discrete QNM (or ringdown) frequencies Leaver (1985). For gravitational perturbations of Kerr, QNM frequencies are labeled by three integers: $\ell$, $m$ and $n$, where $\ell\geq 2$, $|m|\leq\ell$ and $n\geq 0$ Berti et al. (2006a). Typically, $\ell$ is referred to as a polar index, $m$ an azimuthal index, and $n$ an “overtone” index. Respectively, $\ell$, $m$ and $n$ originate from the polar, azimuthal and radial dimensions, as defined in a BH centered frame Leaver (1985).
For physically relevant scenarios, i.e. where the QNM frequencies depend on $\ell$, the spheroidal harmonics have recently been shown to not only be biorthogonal [^2] Two sequences, $\{a_{i}\}_{i=0}^{\infty}$ and $\{b_{j}\}_{j=0}^{\infty}$, are biorthogonal if ${\langle{a_{i}}\,|\,{b_{j}}\rangle}=\delta_{ij}$ Brauer (1964); Christensen (2003). with the so-called adjoint-spheroidal harmonics, but also complete London (2023). This conclusion is based on the fact that $\ell$ is directly related to the eigenvalues of e.g. Jacobi polynomials Fackerell and Crossman (2008). As a result, the spheroidal harmonics may be thought of as a basis of special functions, capable of representing gravitational waves from arbitrary sources.
This is well known to be true of the spherical harmonics, which are a special case of the spheroidal harmonics. Consequently, spherical harmonics play a variety of key roles in gravitational wave theory and data analysis, from the development of perturbative approximations (e.g. Post-Newtonian theory and gravitational wave self force), and the storage of simulated gravitational waves (Numerical Relativity), to the evaluation of gravitational wave signal models for use in parameter inference Newman and Penrose (1966); Goldberg et al. (1967); Blanchet (2014); Abbott et al. (2016d); O’Sullivan and Hughes (2014); Ruiz et al. (2008); Breuer et al. (1977). Analogous applications for the spheroidal and adjoint-spheroidal harmonics have been proposed London and Hughes (2022). Perhaps surprisingly, there has been comparatively little research into whether the QNM’s radial functions are special functions of potentially broad use in gravitational wave theory Green et al. (2023); Lagos and Hui (2022); Yang et al. (2012); Leaver (1986).
In $1973$, Teukolsky derived the differential equation that the radial functions must satisfy, and at the time noted that his radial equation has two regular singular points and one irregular one, meaning that it cannot be transformed into a hypergeometric-type equation for which there are many well known analytic methods of solutions Teukolsky (1973). This appears to have informed a branch of research which seeks to solve Teukolsky’s radial equation using numerical methods, e.g. O’Sullivan and Hughes (2014); Hughes (2000a); Taracchini et al. (2014); Hughes et al. (2019); Islam et al. (2022); Rifat et al. (2019); Campanelli et al. (2001); Krivan et al. (1996); Vishal et al. (2023).
However, in 1985 Leaver presented what was perhaps the first Frobenius-type series solution for the QNM’s radial functions Leaver (1985, 1986). Leaver’s work on QNMs intersects with topics in gravitational wave self-force, in which asymptotic solutions to Teukolskys radial equation play a key role Isoyama et al. (2022); Casals and Zimmerman (2019); Sasaki and Tagoshi (2003); Fujita and Tagoshi (2004).
In parallel, there has been broad and growing interest in a generalization of the hypergeometric equation called the Heun equation Ronveaux and Arscott (1995); Chen and Jing (2023); Fiziev (2009a); Zarrinkamar and Hortaçsu (2018); Dariescu et al. (2021); Magnus et al. (2021); Bonelli et al. (2022). Unlike the hypergeometric equation, which has three regular singular points, the Heun equation has four Ronveaux and Arscott (1995). It was perhaps first pointed out by Leaver and then Fiziev that Teukolsky’s radial equation is an instance of the confluent Heun equation, also known as a generalized spheroidal equation Fiziev (2009a). Fiziev’s work appears to coincide with what is perhaps a recent and multi-disciplinary interest in Heun equationsZarrinkamar and Hortaçsu (2018). For example, it has very recently been observed that a careful consideration of the confluent Heun function’s asymptotic behavior allow them to be applied directly to astrophysically relevant scenarios in particle perturbation theory (e.g. gravitational waves from extreme mass ratio inspirals) Chen and Jing (2023).
Together, methods for evaluating the QNM’s radial functions that are derived from Teukolsky, MST and Heun underlay a vast (and growing) body of research. However, thus far, this work appears to have not fully addressed the following questions:
  1. (Q1)
    The properties of many special functions in physics may be framed using ideas from Sturm-Liouville theory and functional analysis Arfken et al. (2013); Kristensson (2010); Courant and Hilbert (1989); Adkins and Weintraub (2012). There, solutions to differential equations are conceptualized as eigenvectors, where the vector scalar products (i.e. symmetric nondegenerate bilinear forms Adkins and Weintraub (2012)) are defined by an integral whose integrand is weighted such that the differential equations is formally self-adjoint. For example, the scalar product for the QNM’s angular functions can be framed as that of the Jacobi polynomials (i.e. with the Jacobi weight function), and the Jacobi differential equation is self-adjoint with respect to that scalar product Fackerell and Crossman (2008). Further, the Jacobi polynomials are compatible with the spheroidal harmonic’s analytic (i.e. regular and differentiable) structure Fackerell and Crossman (2008). In analogy:
    Is there an appropriate scalar product for the QNM’s radial functions?
  2. (Q2)
    Classical polynomials (i.e. those of Hermite, Laguerre and Jacobi) are known to be solutions to special cases of the hypergeometric equation Abramowitz and Stegun (1964). It is well known that the QNM angular functions are underpinned by spin-weighted spherical harmonics, which are themselves underpinned by Jacobi polynomials Fackerell and Crossman (2008). Therefore, there is good reason to wonder whether an analogous conclusion be developed for the QNM’s radial functions:
    Are the QNM’s radial functions underpinned by classical polynomials, or some generalization thereof? Equivalently, is the overtone label, i.e. $n$, directly related to the order of a complete and orthonormal polynomial sequence?
Regarding (Q1), many studies have encountered the need to perform integrals over solutions to Teukolsky’s radial equation, e.g. Green et al. (2023); Motl and Neitzke (2003); Hughes (2000a); Leaver (1986); Andersson (1997); Jaramillo et al. (2021). Perhaps most relevant to the present article, Leaver (Ref. Leaver (1986)) studied Schwarzschild (i.e. non-spinning) BHs, and used a Green’s function approach to derive various spectral-type expansions, including those for QNMs. Similar approaches have recently been used to study non-linear (particularly quadratic) QNMs of Schwarzschild, e.g. Lagos and Hui (2022). This branch of research intersects with (Q1), however it does not directly draw from Sturm-Liouville theory due to apparent divergences in QNM-type problems that are not present in classical resonance problems [^3] Similarly, Ref. Nollert and Price (1999a) implies that the QNM problem may not be compatible with Sturm-Liouville theory because QNMs damp in the time domain. In Sec. IV of this article, this is shown to not be the case. In paricular, the QNM problem is non-Hermitian; however, this does not preclude the self-adjoint perspective taken in this work. Leaver (1986). Interestingly, Ref. Leaver (1986) observes that exactly such divergence may be overcome by allowing for a radial integration path that is complex valued. In this article, the same idea is encountered for Kerr, and is motivated by coordinate dependence.
In an adjacent branch of research, many studies have made use of what’s called the “formal adjoint” of operators related to QNMs where, in effect, a weight function of $1$ is considered, e.g. Teukolsky and Press (1974); Wald (1979); Green et al. (2023); Dias et al. (2022); Fiziev (2009a); Jaramillo et al. (2021). From that perspective, Teukolsky’s radial equation is not self-adjoint due to the weight function or scalar product used. Consequently, it appears that an investigation of the QNMs’ radial functions in the context of Sturm-Liouville theory with a physically motivated weight has not yet been pursued.
Regarding (Q2), a number studies have attempted to use classical polynomials to represent solutions to the QNMs’ radial functions, e.g. Chung et al. (2023a, b); Ghojogh et al. (2023). These efforts encounter the following mathematical difficulty: while Teukolsky’s radial equation may be similarity transformed to be similar to the equation that governs e.g. Jacobi polynomials, the related transformations differ from those required by the physical boundary conditions Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). In turn, this means that unlike Teukolsky’s angular eigenvalue problem for QNMs, Teukolsky’s radial eigenvalue problem cannot be exactly represented in terms of a standard spectral problem involving classical polynomials. Instead, one encounters a “generalized spectral problem” Chung et al. (2023a); Ghojogh et al. (2023). Although it has been recognized that Teukolsky’s radial equation is a confluent Heun equation, it appears that confluent Heun polynomials have not been used to frame the QNM radial problem as purely spectral until the present work.
In 2014, Cook and Zalutskiy studied the connection between confluent Heun polynomails, “algebraically special” QNMs of nonspinning BH’s, and “total transmission” QNMs in the presence of BH spin Cook and Zalutskiy (2014). There, it was found that for special QNM frequencies, the associated radial functions can be directly identified with select confluent Heun polynomials. With (Q2) in mind, there is good reason to ask whether the confluent Heun polynomials are only relevant to the total transmission QNMs, or whether there exists a single formulation that connects confluent Heun polynomials to all QNM solutions?
This article provides a first-principles development of a radial scalar product for QNMs, thereby affirmatively answering (Q1). Towards (Q2), this article applies the scalar product to confluent Heun polynomials, where it is shown that the polynomials may be defined for any wave frequency, including the complex valued QNM frequencies. While a full treatment of (Q2) is determined to be outside of this article’s scope, significant progress is provided towards an affirmative answer. In particular, this article concludes by showing that Teukolsky’s radial equation is, at least in principle, exactly tridiagonalizable. To this end, the existence of potentially new “canonical confluent Heun polynomials” is conjectured.
II Scope and organization
II.1 Scope
This article discusses QNM solutions to Teukolsky radial equation in the context of isolated Kerr BHs, absent any effects beyond GR. Some aspects of the discussion make minimal physical assumptions, while others are specific to gravitational wave ringdown. Broadly, this work studies the discrete spectra of Teukolsky’s radial equation. By construction, all versions of this equation are predicated on a single QNM frequency (i.e. the radial equation for QNMs may only be defined for a single frequency at a time).
The underlying perspective of this work is that the structure QNM frequencies is influenced by the underlying structure of the discrete spectra of the radial equation for single frequencies. This concept is outlined in Sec. III.1, and it is encapsulated by the overtone number, $n$. It is known that many references refer to the overtone number as being heuristically related to an “apparent” order in the imaginary part of QNM frequencies Berti et al. (2006a); Nollert (1999); Nollert and Price (1999b). That apparent order is equally present in the QNM separation constants, and in the context of this work, each QNM separation constant is one of many radial eigenvalues.
Most mathematical arguments make no assumption about spin weight, $s$. However, $s\in\{-2,+2\}$ are given particular attention due to their correspondence with gravitational radiation. Further, Teukolsky’s radial equation, even when formatted for the QNMs, may be considered for any frequency. Therefore, much of this article’s content is not specific to the QNM’s frequencies, but rather the structure of Teukolsky’s radial equation when formatted with QNM boundary conditions. The reader is cautioned that the author does not aspire to the language of formal mathematical rigor (e.g. the text will not include theorems and lemmas).
This article’s primary focus is a frequency dependent scalar product that is uniquely determined by a choice of radial coordinate and the QNM boundary conditions. While it is noted in Sec. IV.2 that this product is related to a range of ostensibly frequency independent products, a key feature of all such products is that they are conditioned on the frequency parameter used to separate Teukolsky’s equation. In that sense, all such products are frequency dependent: some place the frequency dependence in a pre-factor in front of the QNM radial functions, while others, including the one studies here, place the frequency dependence within the scalar product’s weight function. This concept is expanded upon in Sec. IV.2.
Throughout this article, QNM frequencies are used to evaluate numerical examples due to their relevance to gravitational wave theory and experiment. QNM frequencies for this article have been generated using Leaver’s method via the publicly available positive and qnm Python packages London et al. (2020); Stein (2019). All numerical results presented have been computed with $64$ significant figures of precision via the mpmath package mpmath development team (2023).
II.2 Organization
Section III provides an overview of Kerr QNMs and Teukolsky’s radial equation. Section IV discusses a simplified version of Teukolsky’s radial equation in the context of Sturm-Liouville theory, and uses ideas therefrom to derive a scalar product for the Kerr QNMs. Section V focuses on two complementary numerical methods (direct integration and analytic continuation) for evaluating the scalar product. Section VI presents the limitations of direct integration and analytic continuation, and then compares and contrasts them.
Section VII presents two pedagogical applications of the scalar product. There, a pedagogical construction of the confluent Heun polynomials is provided, along with a derivation their most relevant properties. Section VIII discusses the implications of the scalar product and confluent Heun polynomials on Teukolsky’s radial equation. There, a class of “canonical confluent Heun polynomials” is proposed, and used to argue that Teukolsky’s radial equation may be exactly tridiagonalizable.
Finally, in Sec. IX, further implications of the scalar product and related results are discussed in the context of future work.
III Teukolsky’s equation with QNM boundary conditions
In this section we connect the observable gravitational wave polarizations with this article’s topical starting point, Teukolsky’s radial equation with QNM boundary conditions. Along the way, we will encounter spin weighted fields Newman and Penrose (1966, 1962), the definition of QNMs Nollert (1999); Berti et al. (2016), the importance of a scalar product and polynomials to Teukolsky’s angular equation, and how QNM solutions to Teukolsky’s radial equation may be framed in close analogy with the angular case Leaver (1985). These ideas converge on Teukolsky’s radial equation in a simplified form Cook (2020); Fiziev (2010); Chen and Jing (2023). That equation sets the stage for our application of Sturm-Liouville theory Pinchover and Rubinstein (2005); Abramowitz and Stegun (1964); Arfken et al. (2013); Courant and Hilbert (1989), and the development of a radial scalar product in Sec. IV.
III.1 From gravitational wave strain to the perturbative wave quantity
Any gravitational wave detector will be sensitive to the two strain polarizations predicted by GR, $h_{+}$ and $h_{\times}$ Ruiz et al. (2008); Thorne (1980). The study of gravitational wave plane waves motivates an extremely useful complex valued strain given by
$h\;$ $=\;h_{+}\,-\,i\,h_{\times}\;$ (1a)
$=\;\int_{-\infty}^{t}\int_{-\infty}^{t^{\prime}}\psi_{4}(t^{\prime\prime},r,\theta,\phi)\;dt^{\prime\prime}dt^{\prime}\;.$ (1b)
In Eq. (1a), $h$ is the complex valued gravitational wave strain Ruiz et al. (2008). We will consider it to depend on Boyer–Lindquist coordinates, where time, radius, polar angle and azimuthal angle are denotes by $t$, $r$, $\theta$ and $\phi$ respectively Boyer and Lindquist (2004). In Eq. (1b), $\psi_{4}$ is the $5$’th Weyl scalar in Newman-Penrose notation Ruiz et al. (2008); Newman and Penrose (1962). In that formalism, $\psi_{4}$ is a complex valued scalar field of spin weight s=2s=-2 Newman and Penrose (1966); Teukolsky (1973). Henceforth, complex valued qualities encode or modify the relative importance of $h_{+}$ and $h_{\times}$ within $h$. Within the study of gravitationally perturbed isolated BHs, two Weyl scalars are of particular interest: $\psi_{4}$ (s=2s=-2), and the closely related $\psi_{0}$ ($s=+2$). Both encode information about gravitational waves Teukolsky (1973).
For linear perturbations of Kerr, $\psi_{4}$ is treated in terms of a generic wave quantity, $\psi$:
$\psi\;$ $=\;-(r-ia\cos\theta)^{-4}\;\psi_{4}\;.$ (2)
In Eq. (2), the quantity scaling $\psi$ is a spin coefficient Teukolsky (1973), where $a=|J|/M$ is the BH spin magnitude. We will hold that $G=c=1$, leaving $a$ to have units of the black hole mass, $M$.
The use of $\psi$ is that it satisfies exactly the same partial differential equation as other physically relevant wave quantities. In the case of the first Weyl scalar, $\psi_{0}$, we simply have that $\psi=\psi_{0}$ Newman and Penrose (1962). Whether $\psi$ refers to $\psi_{0}$ or $\psi_{4}$ is toggled by the spin weight of the field, $s$. Thus for $\psi_{0}$ ($s=+2$) or $\psi_{4}$ (s=-2), $\psi$ satisfies Teukolsky’s equation, which is one way of representing Einstein’s equations linearized about the Kerr solution Hughes (2000b); Teukolsky (1973, 1972); Teukolsky and Press (1974); Press and Teukolsky (1973); Mino et al. (1997).
III.2 Teukolsky’s equation and its QNM solutions
If we denote the partial differential operator for Teukolsky’s equation as $\mathcal{L}_{tr\theta\phi}$, with
$\mathcal{L}_{tr\theta\phi}\;=\;\mathcal{L}_{tr\theta\phi}(s,r,\theta,\partial_{t},\partial_{r},\partial_{\theta},\partial_{\phi}),$ (3)
then the schematic form of Teukolsky’s equation is
$\mathcal{L}_{tr\theta\phi}\;\;\psi(t,r,\theta,\phi)\;=\;4\pi\Sigma T\;.$ (4)
In Eq. (3), $\{\partial_{t},\partial_{r},\partial_{\theta},\partial_{\phi}\}$ denote differentiation with respect to the the Boyer-Lindquist coordinates Boyer and Lindquist (2004). An explicit definition of $\mathcal{L}_{tr\theta\phi}$ is provided in Appx. (A). In Eq. (4), $\Sigma$ is the usual quantity that appears in the Kerr metric,
$\Sigma\equiv r^{2}+a^{2}\cos^{2}\theta\;,$ (5)
and $T$ is a source term derived from the stress-energy tensor. Since we are interested in isolated BHs, we have that
$T(t,r,\theta,\phi)\;=\;0\;.$ (6)
The structure of $\mathcal{L}_{tr\theta\phi}$ (see Eq. 3) is known to be such that $\psi$ of the form
$\psi\;\propto\;R(r)\,S(\theta)\,e^{-i\tilde{\omega}{}t}\,e^{-im\phi}\;,$ (7)
separates [^4] The ansatz provided in Eq. (7) obscures the fact that our net solution will depend on the tortoise coordinate via the retarded time Press and Teukolsky (1973); Teukolsky (1973). The radial part of the retarded time is ultimately encoded within $R(r)$. $\mathcal{L}_{tr\theta\phi}$ into two coupled equations: one for $R(r)$ and another for $S(\theta)$. In Eq. (7), $\tilde{\omega}$ is generally a complex valued frequency:
$\tilde{\omega}\;=\;\omega\,-\,i/\tau\;.$ (8)
In Eq. (8), one should think of any physical situation in which an isolated BH experiences a momentary and small perturbation. The post-merger of two sufficiently compact objects is an example of particular importance to astrophysics Helfer et al. (2017); Abbott et al. (2021b, a); Baker et al. (2006). For the simplest scenarios, the resulting gravitational radiation will, qualitatively, ring down Nollert (1999); Press (1971); Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999a). The central frequency of that ringing will be $\omega$, and the exponential decay rate of that ringing will be $1/\tau$, where $\tau$ is strictly positive; otherwise, Eq. (7) would diverge at late times, and the BH would be linearly unstable Vishveshwara (1970); Press and Teukolsky (1973); Press (1971); Whiting (1989).
In general, gravitational radiation will be a sum of such mode-like terms (Eq. 7) weighted by constant factors that depend on spacetime geometry and the perturber Berti and Cardoso (2006); Kamaretsos et al. (2012); London et al. (2014); London (2020). The net spatial dependence of each term in the sum defines a single QNM Berti et al. (2007); Kokkotas and Schmidt (1999); Nollert and Price (1999a). There will always be a period in time for which QNM radiation dominates over other solutions to the homogeneous radial equation, such as power-law tails, prompt emission, and nonlinear QNMs Leaver (1986); Lagos and Hui (2022); Andersson (1997). This inherently poorly defined region of QNM dominance is colloquially referred to as ringdown Press (1971); Nollert (1999).
Properties of QNM solutions can be determined by applying the solution ansatz defined in Eq. (7), and then studying the resulting coupled eigenvalue problems. These are separate radial and angular differential equations, connected by the separation constant, $A$:
$\mathcal{L}{{}_{u}}\,S(u)\;$ $=\;-A\,S(u)$ (9a)
$\mathcal{L}{{}_{r}}\,R(r)\;$ $=\;+A\,R(r)\;.$ (9b)
Only the schematic structure of Eqs. (9a) and (9b) are important here, but reader may find full expressions for $\mathcal{L}{{}_{u}}$ and $\mathcal{L}{{}_{r}}$ in Appx. (A). Henceforth, we will refer to Eq. (9a) as the angular equation, and Eq. (9b) as the radial equation. Respectively, angular problem and radial problem will refer to the solving of each eigenvalue equation.
In Eq. (9), $\cos(\theta), u\mathcal{L}{{}_{u}}$ will be referred to as the angular differential operator, and $\mathcal{L}{{}_{r}}$ will be referred to as the radial differential operator. The angular functions, $S(u)$, are called spheroidal harmonics, and we will refer to $R(r)$ as Teukolsky’s radial functions. Both are known to be closely related to the confluent Heun functions [^5] As noted by Leaver, they are also called generalized spheroidal wave functions Leaver (1986). Fiziev (2010); Cook and Zalutskiy (2014); Cho et al. (2009); Chen and Jing (2023).
III.3 Lessons from Teukolsky’s angular problem: the roles of the scalar product and Jacobi polynomials
Although both angular and radial equations are known to be similar in form (e.g. Ref. Leaver (1986)), there are good reasons to first study the angular equation, and then apply lessons therefrom to the radial equation. Relative to the radial equation, the angular equation is well known to be simpler; as a result, there are many well known and highly accurate methods for computing its global666i.e. series solutions that apply across the physical domain analytic solutions Leaver (1985); Cook and Zalutskiy (2014); O’Sullivan and Hughes (2014); Berti et al. (2006b); Yang et al. (2012). Pertinent to this article is the spectral method presented in Ref. Fackerell and Crossman (2008), and whether two key concepts therein may be applied to the radial problem.
For the convenience of the reader, a modern and self-contained presentation this method is provided in Appx. (B). There, the two key concepts are:
  • (i)
    The structure of the differential equation allows the development of a scalar product that is natural for the solution space. This concept motivates (Q1).
  • (ii)
    The structure of the differential equation points the way to problem specific special functions, namely the Jacobi polynomials. In turn, the Jacobi polynomials may be used to exactly represent analytic solutions. This concept motivates (Q2).
If applicable to the radial problem, concept (i) may enable greater understanding of QNM (spatial) orthogonality, bi-orthogonality, and completeness London (2023). Similarly, concept (ii) may yield a representation of the radial functions that may not only be computationally efficient, but may also lend clear insight into the mathematical meaning of the so-called overtone label, $n$. For these reasons, concepts (i) and (ii) will guide forthcoming discussion of the radial equation.
III.4 QNM solutions of Teukolsky’s radial equation
The study of BH QNMs is traditionally contained within the study of spacetime exterior to the event horizon Regge and Wheeler (1957); Vishveshwara (1970); Press and Teukolsky (1973). In this context, QNM boundary conditions are defined on two asymptotic surfaces: the event horizon and spatial infinity Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). This makes it computationally and algebraically useful to consider a compactified radial coordinate Leaver (1985, 1986),
$\xi\;$ $=\;(r-r_{+})/(r-r_{-})\;.$ (10)
In Eq. (10), $r_{+}=M+\sqrt{M^{2}-a^{2}}$ is the outer Kerr horizon (i.e. the event horizon), and $r_{-}=M-\sqrt{M^{2}-a^{2}}$ is the inner Kerr horizon. The compactified coordinate, $\xi$, has the effect of mapping $r=r_{+}$ to $\xi=0$, $r=r_{-}$ to $\xi\rightarrow$-$\infty$, and $r\rightarrow\infty$ to $\xi=1$.
The QNM boundary conditions are that the gravitational radiation is purely ingoing (towards the BH) at $\xi=0$ and purely outgoing near $\xi=1$. A derivation of these conditions is provided in e.g. Refs.Press and Teukolsky (1973); Cook and Zalutskiy (2014). Since these conditions are imposed between $r=r_{+}$ and $r\rightarrow\infty$, we will only consider $r\in[r_{+},\infty)$, and so $\xi\in[0,1)$.
The QNM boundary conditions are asymptotic in the sense that they are restrictions on the functional form of $\psi$ near the event horizon and spatial infinity. As used by Leaver, these boundary conditions amount to a similarity transformation on the radial problem Ronveaux and Arscott (1995); Leaver (1985, 1986). This transformation proceeds by seeking solutions to the radial problem of the form
$R(r(\xi))\;$ $=\;\mu(\xi)f(\xi)\;.$ (11)
In Eq. (11), $\mu(\xi)$ is a product of functional forms required by the asymptotic boundary conditions,
$\mu(\xi)\;$ $=\;e^{\frac{2i\delta\tilde{\omega}}{1-\xi}}(1-\xi)^{1+2(s-iM\tilde{\omega})}$ (12)
$\quad\quad\times\xi^{-(iM\tilde{\omega}+s)+\frac{i\left(am-2M^{2}\tilde{\omega}\right)}{2\delta}}\;.$
The factors that make up $\mu$ are determined by structure of Frobenius (series) solutions to Eq. (9b). In Eq. (12), the argument of the exponential, as well as the powers of $\xi$ and $1-\xi$ are selected by the physical boundary conditions Cook and Zalutskiy (2014); Fiziev (2010); Leaver (1986). In turn, Eq. (11) strictly requires that $f(\xi)$ be non-zero at $\xi=0$ and $\xi=1$; otherwise the scalings imparted by $\mu$ (i.e. the QNM boundary conditions) would be spoiled. This requirement makes $f(\xi)$ the natural quantity for which to seek a power series solution.
The application of Eq. (11) to radial eigenproblem (Eq. 9b) has the effect of transforming $\mathcal{L}_{r}$ into a new differential operator, $\mathcal{L}_{\xi}$,
$\left[\mu(\xi)^{-1}\,\mathcal{L}{{}_{r}}\,\mu(\xi(r))\right]f(\xi)\;$ $=\;\left[\mu(\xi)^{-1}A\,\mu(\xi)\right]f(\xi)\;$ (13a)
$\mathcal{L}{{}_{\xi}}\,f(\xi)\;$ $=\;A\,f(\xi)\;.$ (13b)
In Eq. (13a), we have applied the solution ansatz given by Eq. (11) to the radial eigenproblem, Eq. (9b). We have then divided by $\mu$. Since $\mu$ is closely related to the singular exponents used in Frobenius’ method, the division is exact Leaver (1985). The result, Eq. (13b), is a transformed eigenproblem that has the same eigenvalues, $A$. It is in this sense that $\mathcal{L}_{\xi}$ is formally similar to $\mathcal{L}_{r}$ Axler (2015); Lax (2002); Christensen (2003),
$\mathcal{L}{{}_{\xi}}\;$ $=\;\mu(\xi)^{-1}\,\mathcal{L}{{}_{r}}\,\mu(\xi(r))\;.$ (14)
While we refer to the effect of Eqs. (13) and (14) as a similarity transformation, it is sometimes referred to as an elementary power transformation, or a homotopic transformation Ronveaux and Arscott (1995); Chen and Jing (2023); Cho (2020). The reader should note that Eqs. (13a) and (14) contain a coordinate transformation ($r$ to $\xi$), as well as a similarity transformation.
In Eq. (14), the new differential operator, $\mathcal{L}_{\xi}$, has the form given by Eq. (15).
$\mathcal{L}{{}_{\xi}}\;$ $=\;(\text{C}_{0}+\text{C}_{1}(1-\xi))+\left(\text{C}_{2}+\text{C}_{3}(1-\xi)+\text{C}_{4}(1-\xi)^{2}\right)\partial_{\xi}+\xi(\xi-1)^{2}\partial_{\xi}^{2}$ (15)
There, the constants, $\text{C}_{0}$ to $\text{C}_{4}$, are defined in Eq. (16) using the physical parameters discussed previously.
$\delta\;$ $=\;\sqrt{M^{2}-a^{2}}$ (16a)
$\text{C}_{0}\;$ $=\;-2am\tilde{\omega}-2i\tilde{\omega}(-\delta+M(2s+1))$ (16b)
$\quad\;\;+\tilde{\omega}^{2}(\delta+M)(\delta+7M)$
$\text{C}_{1}\;$ = $=\;8M^{2}\tilde{\omega}^{2}+s(4iM\tilde{\omega}-1)+6iM\tilde{\omega}-1$ (16c)
$\quad\;\;-(4M\tilde{\omega}+i)\frac{\left(am-2M^{2}\tilde{\omega}\right)}{\delta}\;$
$\text{C}_{2}\;$ = $=\;4i\delta\tilde{\omega}$ (16d)
$\text{C}_{3}\;$ $=\;-2(s+1)+4i\tilde{\omega}(M-\delta)$ (16e)
$\text{C}_{4}\;$ $=\;s+3-6iM\tilde{\omega}+\frac{i\left(am-2M^{2}\tilde{\omega}\right)}{\delta}$ (16f)
Equation (15) has a few special properties. We have started with Boyer-Lindquist coordinates, but one also arrives at Eq. (15) if one instead begins with Kerr Ingoing or Kerr Outgoing coordinates Teukolsky and Press (1974). One also arrives at Eq. (15) (albeit with $s\rightarrow-s$ and $A\rightarrow A-2s$) if one uses the alternative tetrad choice discussed in Ref. Teukolsky and Press (1974), which amounts to $t\rightarrow-t$ and $\phi\rightarrow-\phi$. As noted in e.g. Refs. Teukolsky and Press (1974); Whiting (1989), the use of either $s=2$ or s=2s=-2 results in mathematically different descriptions of the same gravitational radiation.
Equation (15) has regular singular points at $\xi\in\{-\infty,0\}$ and one irregular singular point at $\xi=1$. Thus Eq. (15) is an instance of the confluent Heun differential operator Ronveaux and Arscott (1995). This is also true in the Schwarzschild limit. While it is sometimes said that Ref. Leaver (1986) treats Eq. (15) as what is called a generalized spheroidal operator, we note here that this is a slight misnomer Ronveaux and Arscott (1995); Zarrinkamar and Hortaçsu (2018). Equation (15) may only be transformed into a generalized spheroidal operator if one applies an asymptotic boundary condition at $r=r_{-}$ that is inconsistent with QNMs Leaver (1985); Cook and Zalutskiy (2014). Appendix (C) provides an expanded disambiguation between different forms of the confluent Heun equation and the radial equation.
Equations (13b-16) are commonly used to define series solutions for $f(\xi)$. The requirement that these series solutions converge on $\xi\in[0,1)$ happens to be equivalent to finding eigenvalues, $A$, such that certain continued fractions are zero Leaver (1985, 1986); Mano et al. (1996). In particular, see Ref.Leaver (1986) for a proof of convergence for series solutions, example applications of continued fractions to the radial and angular eigen-problems, and how the eigenvalues may be determined by searching for roots of continued fractions.
From the use of continued fractions, values of $A$ are well known to be discrete. However, the use of continued fractions has arguably obscured deeper understanding of the distribution of eigenvalues, $A$, and their related radial (labeled in $n$) and angular functions (labeled in $\ell$ and $m$). In search of that understanding, it appears that there has been relatively little application of Sturm-Liouville theory to Eq. (15) in the BH context Leaver (1985); Nollert (1999); Berti and Cardoso (2006); Ripley (2022); Lin and Qian (2023); Press (1971); Motl and Neitzke (2003); Ronveaux and Arscott (1995); Zimmerman et al. (2015). Nevertheless, since the Kerr QNMs’ eigenvalues are well known to be unique for most physical scenarios  [^7] There is some ambiguity in how one defines the eigenvalues. In particular, the fact that $\text{C}_{0}$ is non-zero means that the eigenvalue that one starts with, $A$, need not strictly define the eigenvalue that is used in Eq. (9a). Thus in e.g. the limit that $a=0$, where $A=\ell(\ell+1)-s(s+1)$ for every radial eigenfunction, the effective eigenvalue, $A-\text{C}_{0}$, may be a unique function of $\tilde{\omega}$.  (e.g.  London and Fauchon-Jones (2019); Cook (2020); Seidel (1989); Yang et al. (2012)), the QNM’s themselves are necessarily linearly independent [^8] More precisely, finite sets of QNMs will be linearly independent when their eigenvalues are unique. Infinite sets of QNMs will be minimal if their eigenvalues remain distinct as some order label on the set goes to infinity Christensen (2003); London (2023)., and thus well suited to Sturm-Liouville theory Courant and Hilbert (1989).
IV Definition of a radial scalar product
The aim of Sturm-Liouville theory is to understand the existence and asymptotic behavior of eigenvalues and eigenfunctions Pinchover and Rubinstein (2005); Arfken et al. (2013); Courant and Hilbert (1989). Here, we will focus on a facet of the theory that pertains to whether there exists a scalar product and an associated weight function, for which $\mathcal{L}_{\xi}$ is self-adjoint. While a scalar product may be developed under which the problem is self-adjoint (symmetric under scalar products), it is still the case the the QNM radial problem is non-Hermitian; e.g. it has non-trivially complex eigenvalues.
The immediate aim of what follows is to derive a scalar product and related weight function such that Eq. (15) is formally self-adjoint. In other words, a function, $\text{W}(\xi)$, is determined such that $\text{W}(\xi)\mathcal{L}_{\xi}$ is in Sturm-Liouville form. As the scalar product will ultimately be an integral over $\xi$ involving $\text{W}(\xi)$, the behavior $\text{W}(\xi)$ at the domain boundaries is addressed in Sec. V.
IV.1 The scalar product and the weight function
By scalar product, we mean a symmetric bilinear form that takes two non-singular functions [^9] The reader should not confuse the non-singular but otherwise arbitrary function, $\mathrm{a}(\xi)$, with the BH spin, $a$., $\mathrm{a}(\xi)$ and $\mathrm{b}(\xi)$, and performs an operation on them that results in a number Adkins and Weintraub (2012); Gohberg et al. (2006). In less technical parlance, we simply mean a bra-ket operation that is similar to what one encounters in quantum mechanics:
${\langle{\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;=\;\int_{0}^{1}\,\mathrm{a}(\xi)\mathrm{b}(\xi)\;\text{W}(\xi)\;d\xi\;.$ (17)
In Eq. (17), $\text{W}(\xi)$ is a called a measure, or weight function. It will soon be the focus of our discussion.
Since there is no conjugation involved in Eq. (17), the scalar-product shown is not conjugate-symmetric. This feature distinguishes it from an inner-product, and allows analysis of $\mathcal{L}_{\xi}$, irrespective of its complex valued coefficients. The analysis of immediate interest pertains whether $\text{W}(\xi)$ may be defined such that $\mathcal{L}_{\xi}$ is formally self-adjoint Pinchover and Rubinstein (2005); Courant and Hilbert (1989).
The adjoint of $\mathcal{L}_{\xi}$ is defined to be an operator, $\mathcal{L}_{\xi}$$^{\dagger}$, such that
${\langle{\mathrm{a}}\,|\,{\mathcal{L}{{}_{\xi}}\mathrm{b}}\rangle}\;$ $=\;{\langle{{\mathcal{L}{{}_{\xi}}}^{\dagger}\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;.$ (18)
Chapter $5$ Section 3 of Ref. Courant and Hilbert (1989) provides a detailed overview of how $\mathcal{L}_{\xi}$$^{\dagger}$ may be explicitly defined using integration by parts. In that derivation, the fact that the coefficient of the second derivative term in Eq. (15) vanishes at $\xi=0$ and $\xi=1$ is key to Sturm-Liouville boundary conditions being satisfied Pinchover and Rubinstein (2005). There, it is also shown that $\mathcal{L}_{\xi}$ is formally self-adjoint, i.e.
$\mathcal{L}{{}_{\xi}}\;$ $=\;{\mathcal{L}{{}_{\xi}}}^{\dagger}\;$, (19)
if it also holds that
$\mathcal{L}_{\xi}\;$ $=\;p_{0}(\xi)\;+\;\text{W}(\xi)^{-1}\,\partial_{\xi}\left[\,(\text{W}(\xi)\,p_{2}(\xi))\,\times\,\partial_{\xi}\,\right]$ (20a)
$=\;p_{0}(\xi)\;+\;\frac{\partial_{\xi}(\text{W}(\xi)\,p_{2}(\xi))}{\text{W}(\xi)}\,\partial_{\xi}\;+\;p_{2}(\xi)\partial^{2}_{\xi}\;.$ (20b)
In Eq. (20), $p_{0}(\xi)$ and $p_{2}(\xi)$ are the coefficient functions seen in zero’th and second derivative terms of Eq. (15), respectively. If we define $p_{1}$ to be the coefficient function of $\partial_{\xi}$ in Eq. (15), then
$p_{0}\;$ $=\;\text{C}_{0}+\text{C}_{1}(1-\xi)\;,$ (21a)
$p_{1}\;$ $=\;\text{C}_{2}+\text{C}_{3}(1-\xi)+\text{C}_{4}(1-\xi)^{2}$ (21b)
$p_{2}\;$ $=\;\xi(\xi-1)^{2}\;.$ (21c)
In Eq. (20a), the appearance of the weight function is a direct result of Eqs. (18) and (19). To go from Eq. (20a) to Eq. (20b), we have applied the product rule (a.k.a. Leibniz’s rule), to $\partial_{\xi}\,\left[(\text{W}p_{2})\times\partial_{\xi}\right]$. For $\mathcal{L}_{\xi}$ in Eq. (15) to be equivalent to that in Eq. (20b), it must be the case that coefficients of $\partial_{\xi}$ match,
$p_{1}\;$ $=\;\text{W}^{-1}\,\partial_{\xi}(\text{W}\,p_{2})$ (22a)
$\text{W}^{-1}\,\partial_{\xi}\text{W}\;$ $=\;p_{2}^{-1}(p_{1}-\partial_{\xi}p_{2})\;.$ (22b)
In Eq. (22a) we have simply equated the first derivative coefficient of Eq. (20b) with that of Eq. (15). In Eq. (22b), we have used Leibniz’s rule to expand the right-hand-sid of Eq. (22a), and then we have rearranged terms to emphasize that Eq. (22b) is a linear differential equation whose left-hand-side is equal to $\partial_{\xi}\ln\text{W}$.
Equation (22b) can be solved for W using the coefficient functions defined in Eq. (21),
$\text{W}(\xi)\;$ $=\;\,\frac{1}{p_{2}(\xi)}\,\exp\left(\int^{\xi}\frac{p_{1}(\xi^{\prime})}{p_{2}(\xi^{\prime})}\,d\xi^{\prime}\,\right)$ (23a)
$=\;e^{\frac{\text{C}_{2}}{1-\xi}}\;(1-\xi)^{-\text{C}_{2}-\text{C}_{3}-2}\;\xi^{\text{C}_{2}+\text{C}_{3}+\text{C}_{4}-1}\;.$ (23b)
In Eq. (23a), the integration constant has been chosen to be zero, and in Eq. (23b) we explicitly evaluate the integral using Eqs. (21b) and (21c).
It will soon be useful to simplify our notation for $\mathrm{W}(\xi)$ by defining a shorthand for combinations of parameters,
$\text{B}_{0}\;$ $=\;\text{C}_{2}+\text{C}_{3}+\text{C}_{4}-1$ (24a)
$\text{B}_{1}\;$ $=\;-\text{C}_{2}-\text{C}_{3}-2$ (24b)
$\text{B}_{2}\;$ $=\;\text{C}_{2}\;.$ (24c)
Consequently, the weight function takes a less cumbersome form,
$\text{W}(\xi)\;=\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;,$ (25)
and the scalar product between two any two arbitrary functions, $\text{a}(\xi)$ and $\text{b}(\xi)$, may now be written as,
${\langle{\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;=\;\int_{0}^{1}\,\mathrm{a}(\xi)\,\mathrm{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (26)
IV.2 Properties and non-uniqueness of the weight function
With Eq. (26), we arrive at one of this article’s key results. Equation (26) along with Eqs. (16) and (24) define a scalar product for which Teukolsky’s radial equation, Eq. (15), is self-adjoint. In turn, self-adjointness, in principle, implies that as long as the eigenvalues may be defined, and that they are unique, then the eigenfunctions are orthogonal, and very likely complete Ronveaux and Arscott (1995). However, there are good reasons why the explicit demonstration of orthogonality and completeness are beyond the scope of this article [^10] In particular, one route to demonstrating completeness of the single-frequency radial functions would be to construct an isomorphism between them, and a known basis London (2023); Christensen (2003).. The most pressing reason has to do with the fact that $\text{W}(\xi)$, naively, appears to diverge on the event horizon, and at spatial infinity. Understanding that apparent divergence and its consequences is preliminary to further investigation.
Visual inspection of Eq. (23) points to the origin of these divergences. In Eq. (23a), the factor
$p_{2}(\xi)^{-1}\;=\;\xi^{-1}(1-\xi)^{-2}$ (27)
clearly creates a potential for the weight function to become singular at $\xi=0$ and $\xi=1$. For this reason, the radial eigen-problem may be said to be of the singular [^11] Note that e.g. the Jacobi and Legendre polynomial problems are also singular in this way Arfken et al. (2013). Sturm-Liouville type Pinchover and Rubinstein (2005). In Eq. (23b), we see that the net exponents of $\xi$ and $1-\xi$ may still prevent divergence if their total real part is positive. Similarly, the exponential factor may diverge if the real part of $\text{C}_{2}=4i\delta\tilde{\omega}$ is greater than zero.
Thus, to determine the physical conditions in which $\text{W}(\xi)$ diverges, one might embark on an analysis of the sign of parameters within Eq. (23b) or, equivalently, Eq. (25). While this is a somewhat cumbersome exercise, it is easy to see that the exponential factor,
$e^{\text{C}_{2}/(1-\xi)}\;=\;e^{4i\delta\tilde{\omega}/(1-\xi)}\;,$ (28)
will generally diverge at spatial infinity. For that we need to recall the discussion below Eq. (8): Since the imaginary part of $\tilde{\omega}$ must be negative for time-domain stability, the real part of $4i\delta\tilde{\omega}/(1-\xi)$ is generally positive for the QNMs,
$\text{Re}\,\frac{4i\delta\tilde{\omega}}{1-\xi}\;=\;\frac{4\delta}{\tau(1-\xi)}\;>\;0\;.$ (29)
The inequality in Eq. (29) holds because $\delta=\sqrt{M^{2}-a^{2}}>0$ and $\tau>0$. Thus it would appear that $\text{W}(\xi)$ will always diverge on at least one boundary. This appearance happens to be just that.
Above, we have implicitly assumed that $\xi=0$ and $\xi=1$ are approached from the real line. While that is a physically meaningful assumption, previous studies have found it to be an unnecessary practical impediment Leaver (1986); Green et al. (2023); Motl and Neitzke (2003); Mano et al. (1996). In particular, when evaluating the scalar product, $\xi$ need not be a real valued quantity. This point is explored in the two sections that follow. A kind of “physicist’s analytic continuation” is detailed in Sec. V.1, and then a complementary analytic continuation technique is presented in Sec. V.2.
We conclude this section by pointing out two alternative perspectives on the weight function. The first hinges on the fact that Teukolsky’s equation allows for the definition of a weight function prior to applying the QNM boundary conditions. The second pertains to whether the weight function is unique, once the QNM boundary conditions have been applied.
The first perspective is that the form of the weight function is closely linked to the scaling function, $\mu(\xi)$, defined in Eq. (12). In particular, it is easy show (see e.g. Eq. 145) that
$\text{W}(\xi)\;$ $=\;\mu(\xi)^{2}\,\text{W}_{0}(\xi)\;.$ (30)
In Eq. (31), $\text{W}_{0}$ is the radial equation’s weight function before applying the similarity transformation shown in Eq. (15),
$\text{W}_{0}(\xi)\;$ $=\;\xi^{s}\,(1-\xi)^{-2(s+1)}\;.$ (31)
This weight function is independent of $\tilde{\omega}$, $m$ and $a$, and it is directly relevant for the original radial functions, $R(\xi)$ (see Eq. 9b). In essence, Eq. (30) defines how the weight function changes under similarity transformations on the main problem. One way to arrive at Eq. (30) is to note that such transformations preserve the value of scalar products. While $\text{W}_{0}$ has already been used in operator perturbation theory (e.g. Zimmerman et al. (2015)), it is not directly considered here because the corresponding differential operator does not support polynomial solutions that are compatible with the QNM boundary conditions.
The second perspective is that $\text{W}(\xi)$ is not the only weight function for which $\mathcal{L}_{\xi}$ is formally self-adjoint. It is left as an exercise for the reader to show that one can actually enforce a strictly positive and real valued weight function,
$\mathcal{W}(\xi)\;=\;|\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}|\;.$ (32)
This exercise follows exactly the steps discussed in Eqs. (20-23), but with the requirement that the weight function is wrapped within an absolute value. Further, one must use the fact that $\partial_{\xi}\mathcal{W}=\text{sign}(\text{W})\,\partial_{\xi}\text{W}$. While $\mathcal{W}(\xi)$ is amenable to the direct integration approach that we will discuss shortly, it is not at all clear to the author how it may be used with the analytic continuation approach that will be discussed in Sec. V.2. Further, there do not appear to be significant gains in using $\mathcal{W}(\xi)$ rather than $\text{W}(\xi)$. Since complexification of the radial coordinate is also needed for $\mathcal{W}(\xi)$, scalar products will in general still be complex valued. $\mathcal{W}(\xi)$ is noted for potential future interest. Henceforth, this article will focus on the evaluation of scalar products with the comparatively simpler weight function, $\text{W}(\xi)$.
Refer to caption Refer to caption Refer to caption
Figure 1: Examples integration paths (white curves) against weight function contours (colored contours) for the $(s,\ell,m,n)=(-2,2,2,3)$ QNM at three BH spins: (left) $a/M=0.7$, $M\tilde{\omega}=0.4713-0.5843i$, (center) $a/M=0.99$, $M\tilde{\omega}=0.8695-0.2058i$, and (right) $a/M=0.999$, $M\tilde{\omega}=0.9558-0.07299i$. Contours for the real and imaginary parts of the weight function, $\text{W}(\xi)=\mathrm{Re}(\text{W})+i\,\mathrm{Im}(\text{W})$, are shown. For $\mathrm{Re}(\text{W})$, colors between cyan and magenta represent dimensionless values between 50-50 and $50$ respectively. For $\mathrm{Im}(\text{W})$, colors between magenta and yellow represent values between 50-50 and $50$ respectively. Convergence of contours at spatial infinity ($\mathrm{Re}(\xi)=1$) corresponds to the divergence of the weight function. Horizontal contours along $\mathrm{Re}(\xi)<0$ and $\mathrm{Re}(\xi)>0$ are branch cuts. For left, center and right panels, coordinate parameters, $(K_{0},K_{2})$, are $(0.8530,23.4621)$, $(0.7105,1.4010\times 10^{-4})$ and $(0.1585,2.5088\times 10^{-4})$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
V Evaluation of scalar products
The scalar product developed in the previous section has been copied below,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (33)
One of the simplest ways to evaluate Eq. (33) is direct integration (i.e. explicit numerical integration via discretization of the domain). However, divergence of the weight function at spatial infinity or the BH horizon means that naive integration along real values of $\xi$ is bound to fail. So other means must be considered.
In Sec. V.1, we make use of coordinate dependence to develop a method to directly integrate Eq. (33). In Sec. V.2, we note that Eq. (33) coincides with the integral representation of the Tricomi confluent hypergeometric function, also known as the Kummer function of the second kind Arfken et al. (2013); Chen and Dai (2010). That fact, along with the well known analytic continuation of e.g. the Gamma function to complex inputs, is used to analytically evaluate Eq. (33). This section concludes with a brief comparison of the pros and cons of direct integration and analytic continuation.
V.1 Direct integration
An extremely useful lesson from calculus is that, given an integral, we may change the coordinate representation of its integrand without changing the value of the integral. So while the integrand of Eq. (33) is coordinate dependent, its value, $\langle{\text{a}}\,|\,{\text{b}}\rangle$}, is coordinate independent. This concept leads us to seek a new real valued coordinate, $z$, such that the integrand of Eq. (33) does not diverge. Henceforth, we will use $\xi(z)$ to refer to a complex valued integration path, and $\xi$ to refer to the quantity defined by Eq. (10).
The following definition of $\xi(z)$ has the desired effect of augmenting singular behavior at the horizon and spatial infinity,
$\xi(z)\;=\;\frac{{\text{x}_{2}}z^{{\text{x}_{0}}}}{({\text{x}_{2}}+1)z^{{\text{x}_{0}}}-1}\;.$ (34)
In Eq. (34), $\text{x}_{0}$ and $\text{x}_{2}$ are coordinate parameters that modify $\text{B}_{0}$ and $\text{B}_{2}$, respectively. We will consider $\text{Re}(x0)>0$, such that $\xi(0)=0$. The dependence of Eq. (34) on $\text{x}_{2}$ is motivated by the complex radial coordinate used in Ref. Mano et al. (1996). It may also be developed by placing the Laguerre polynomial equation into a $\xi$-like compactified coordinate, and then studying that system’s weight function. The dependence of Eq. (34) on $\text{x}_{0}$ is motivated by the observation that $z^{\text{x}_{0}}$ is the simplest apparent coordinate modification that impacts singular behavior arising from $\text{B}_{0}$.
The effect of this coordinate choice is found by applying Eq. (34) to the scalar product’s integrand,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}=\int_{0}^{1}\text{a}(\xi(z))\,\text{b}(\xi(z))\;\left[\text{W}(\xi(z))\;\partial_{z}\xi(z)\right]\;dz\;.$ (35)
In Eq. (35), we have used the fact that $d\xi=\partial_{z}\xi(z)\,dz$ to write the integral explicitly in $z$. According to Eq. (34), there may exist coordinate parameters, $\text{x}_{0}$ and $\text{x}_{2}$, such that $\xi(z=0)=0$, and $\xi(z=1)=1$, which would mean that the limits of integration do not change.
Within Eq. (34), the quantity in square brackets is of particular practical use. It is an effective weight function,
$\text{W}_{\text{eff.}}(z)\;$ $=\;\text{W}(\xi(z))\;\partial_{z}\xi(z)$ (36a)
$=\;e^{\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}}\;z^{\text{x}_{0}(\text{B}_{0}+1)-1}\;\left(1-z^{\text{x}_{0}}\right)^{\text{B}_{1}}\;$ (36b)
$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
While Eq. (34) provides a coordinate that modifies the boundary behavior of $\text{W}(\xi)$, it is Eq. (36) that must be analyzed to determine the special values of $\text{x}_{0}$ and $x_{2}$ that prevent the effective weight function from diverging at the domain boundaries. In essence, we have two independent conditions: (i) $\text{W}_{\text{eff.}}(0)=0$, and (ii) $\text{W}_{\text{eff.}}(1)=0$. We also have two unknowns: $\text{x}_{0}$ and $\text{x}_{1}$. Thus, $\text{x}_{0}$ and $\text{x}_{1}$ may be determined (up to positive constants) by careful analysis of $\text{W}_{\text{eff.}}$.
This analysis is somewhat tedious, but tractable. First, one should focus on the factor of $z^{\text{x}_{0}(\text{B}_{0}+1)-1}$ within $\text{W}_{\text{eff.}}$. If $\text{x}_{0}$ is chosen appropriately, then the real part of $\text{x}_{0}(\text{B}_{0}+1)-1$ should equal some positive constant, $\text{K}_{0}$, thereby allowing $\text{W}_{\text{eff.}}$ to go to zero as $z$ goes to zero. Since, $\text{B}_{0}$ is complex valued, $\text{x}_{0}$ must be also. These ideas result in the following constraints,
$\text{K}_{0}$ $=-1+\left[1+\text{Re}(\text{B}_{0})\right]\text{Re}(\text{x}_{0})-\text{Im}(\text{B}_{0})\text{Im}(\text{x}_{0}),$ (37a)
$\text{K}_{0}$ $>0\;.$ (37b)
If we assume that $\text{Re}(\text{x}_{0})$ only depends on $\text{Re}(\text{B}_{0})$, then then the result for $\text{x}_{0}$ follows,
$\text{x}_{0}\;$ $=\;\text{K}_{0}\,\frac{\text{sgn}(\mathrm{Re}(\text{B}_{0})+1)}{\mathrm{Re}(\text{B}_{0})+1}$ (38)
$\quad\;+\;i\,\frac{-1\,+\,\text{K}_{0}\,[\text{sgn}(\;\mathrm{Re}(\text{B}_{0})+1\;)-1]}{\mathrm{Im}(\text{B}_{0})}\;.$
Note that, in Eq. (38), $\text{Re}(\text{x}_{0})>0$, as is needed for $\xi|_{z=0}$ to be zero.
Given, $\text{x}_{0}$, one may then determine $\text{x}_{2}$. Returning to $\text{W}_{\text{eff.}}$, if its exponential factor has a negative argument as $z\rightarrow 1$, then $\text{W}_{\text{eff.}}$ is assured to be regular in that limit. To that end, one should consider $e^{\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}}$, and seek $\text{x}_{2}$ such that
$\lim_{z\rightarrow 1}\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}\;$ $\sim\;-\frac{\text{K}_{2}}{1-z}\;,$ (39a)
$\text{K}_{2}\;$ $>0\;.$ (39b)
To evaluate the left-hand-side of Eq. (39a), one can use the binomial theorem [^12] For this it is useful to work with $z^{\prime}=1-z$.; this results in a quantity with the same form as the right-hand-side of Eq. (39a). Upgrading the asymptotic equivalence in Eq. (39a) to an equality yields the following solution for $\text{x}_{2}$,
$\text{x}_{2}\;$ $=\;\text{K}_{2}\;\text{x}_{0}\,/\,\text{B}_{2}\;.$ (40)
Note that, since $\text{x}_{0}$ is complex valued, the factor of $\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}$ within $\text{W}_{\text{eff.}}$ will never encounter a root. Thus it is of little practical concern.
In Eqs. (38) and (40) we define the values of the $\text{x}_{0}$ and $\text{x}_{2}$ are allowed to take such that $\text{W}_{\text{eff.}}$ is regular (actually zero) at the event horizon and spatial infinity. In doing so, we have transferred the coordinate dependence to two new, strictly positive, parameters, $\text{K}_{0}$ and $\text{K}_{2}$. What remains is the possibility that $|\xi(z)|\gg 1$.
If $\text{K}_{0}$ and $\text{K}_{2}$ are not chosen such that $|\xi(z)|\lessapprox 1$, then numerical evaluation of Eq. (35) can encounter extremely large numbers. This numerically problematic possibility can be avoided if one determines $\text{K}_{0}$ and $\text{K}_{2}$ by requiring that $\xi(z)$ is contained as much as possible within a unit circle centered about $\xi=1/2$. This may be accomplished via the following minimization,
$\{\text{K}_{0},\text{K}_{2}\}=\text{argmin}_{\text{K}_{0},\text{K}_{2}}S(\text{K}_{0},\text{K}_{2},z)$ (41a)
where
$S(\text{K}_{0},\text{K}_{2},z)\;$ $=\;|1/2-\text{mean}_{z}(|\xi(z)|)|$ (41b)
$\quad\;+\;\text{max}_{z}|\text{Im}\,\xi(z)|\;+\;|\text{min}_{z}\;\text{Re}\,\xi(z)|\;.$
In Eq. (41), argmin outputs the values for which $S$ is minimized in $\text{K}_{0}$ and $\text{K}_{2}$, $\text{mean}_{z}$ computes the unweighted mean over values of $z$, $\text{min}_{z}$ returns the smallest value of a quantity in $z$, and $\text{max}_{z}$ returns the largest value of a quantity in $z$.
In practice, one has a number of remaining options for actually evaluating Eq. (35). For the evaluation of Eq. (41), it is found to be useful to densely sample, on log-scale, $z$ near 0 and $1$. The coordinate freedom associated with $\text{x}_{0}$ introduces oscillations into $\xi(z)$ which must be well resolved for the accurate determination of $\text{K}_{0}$ and $\text{K}_{2}$. Similarly, for the evaluation of Eq. (35), one will want to densely sample near the boundaries; it is found that $z=\frac{1}{2}(1+\sin(t\pi/2))$, with $t$ uniformly spaced on (1,1)(-1,1), is sufficient for this purpose. For the determination of $\text{K}_{0}$ and $\text{K}_{2}$, and for the evaluation of Eq. (35), it is found that at least $10^{4}$ points in $z$ are needed for robust results. With these details in hand, it is now possible to compute the integration path, $\xi(z)$, and related scalar-products, Eq. (35), using e.g. $5^{\text{th}}$ order spline integration Virtanen et al. (2020).
Figure 1 shows three example integration paths, $\xi(z)$, set against contours of the weight function, $\text{W}(\xi)$, in the complex plane. In focus is the effect of BH spin $a$. The result for $a=0$ is qualitatively similar to that for $a=0.7$. Branch cuts are visible as overlapping horizontal contours along the real line for $\xi<0$ and $\xi>1$. The shape of integration paths are seen to vary significantly with increasing BH spin, changing most drastically near the extremal limit, $a=M$.
For $a=0.99$ (the central panel), the integration path appears to cross the branch cut [^13] For complex numbers $y$, We define the principle branch of $\log(y)$ by taking $\arg(y)\in(-\pi,\pi)$ and place the branch cut on the line defined by $\Re(y)<0$. As seen in Fig. 1, this convention results in a branch cut along $Re\xi<0$ (for powers of $\xi$), and another along $Re\xi>1$ (for powers of $1-\xi$). many times. The top panel of Fig. 2 shows this in detail. There, the apparent crossing of branch cuts is an artifact of presentation: $\xi(z)$ actually runs along all branches of the Riemann surface defined by $\text{W}(z)$, but in Fig. 1, only a single branch of $\text{W}(\xi)$ is shown Kuijlaars et al. (2003). This can be understood by noting that the transformed scaler-product, Eq. (35), uses an integration path that follows the real line, where there are no branch cuts. This is shown explicitly in the bottom panel of Fig. 2.
V.2 Analytic continuation
Figure 2: Examples integration paths (white curves) against weight function contours (colored contours) for the $(s,\ell,m,n)=(-2,2,2,3)$ QNM at BH spin $a/M=0.99$. (top) A zoom-in of the central panel of Fig. 1 with axes centered around the BH horizon at $\xi=0$. (bottom) Visualization of the coordinate transformed integration path and effective weight function for the same case. Contour formatting and coordinate parameters are identical to those in Fig. 1.
The scalar product developed in Sec. IV is,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (42)
The method described in the previous section uses coordinate freedom to extend the domain upon which Eq. (42) can be evaluated. This is the basic purpose of the set of techniques known as analytic continuation Ahlfors ; Ronveaux and Arscott (1995). It happens that there is a more direct, and ultimately more robust, route to evaluating Eq. (42) with analytic continuation.
This alternative route begins with the observation that if $\text{a}(\xi)=\text{b}(\xi)=1$, then the left-hand-side of Eq. (42) coincides exactly with the integral representation of the Tricomi [^14] The nomenclature for these functions appears to be particularly inconsistent across sources. Some readers may find Refs. DLMF (a) and Chen and Dai (2010) of particular use. confluent hypergeometric function Arfken et al. (2013); Chen and Dai (2010); Abramowitz and Stegun (1964); DLMF (a); Leaver (1986),
$\int_{0}^{1}\,$ $\xi^{\text{B}_{0}}\,(1-\xi)^{\text{B}_{1}}\,e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi$ (43)
$=\;e^{\text{B}_{2}}\,\Gamma(\text{B}_{0}+1)\,U(1+\text{B}_{0},-\text{B}_{1},-\text{B}_{2})\;.$
In Eq. (43), $\Gamma(x)$ is the usual Euler-Bernoulli Gamma function, and $U(\text{x},\text{y},\text{z})$ is the Tricomi confluent hypergeometric function,
$U(\text{x},\text{y},\text{z})\;=\;$ $\frac{\Gamma(1-\text{y})}{\Gamma(\text{x}-\text{y}+1)}M(\text{x},\text{y},\text{z})$ (44)
$\;+\;\frac{\Gamma(\text{y}-1)}{\Gamma(\text{x})}\text{z}^{1-\text{y}}M(\text{x}-\text{y}+1,2-\text{y},\text{z})\;.$
Both may be evaluated for complex inputs by considering the appropriate integration contour Whittaker and Watson (1996). Finally, in Eq. (44), $M(\text{x},\text{y},\text{z})$ is the confluent hypergeometric function of the first kind (See e.g. Eq. $13.2.42$ of Ref. DLMF (a)).
The Gamma function is constructed to extend the factorial function to the complex numbers. The confluent hypergeometric function is a product of Gamma functions and a power series in z, and so it is straightforward to evaluate.
This parade of special functions has the following use: Every $\text{a}(\xi)$ and $\text{b}(\xi)$ of interest to Eq. (42) will be smooth, and so possesses a series expansion,
$\text{a}(\xi)\;=\;\sum_{j=0}\,\text{a}_{j}\,\xi^{j}\quad\text{ and }\quad\text{b}(\xi)\;=\;\sum_{k=0}\,\text{b}_{k}\,\xi^{k}\;.$ (45)
This allows every scalar product to be expressed as a sum over scalar products of monomials,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\sum_{j,k}\;\text{a}_{j}\,\text{b}_{k}\,\left<\xi^{j+k}\right$>\;. (46)
In Eq. (46), $\langle\xi^{p}\rangle$ is the $p^{\text{th}}$ monomial moment defined as,
$\left<\xi^{p}\right$>\; $=\;\int_{0}^{1}\,\xi^{p}\,\text{W}(\xi)\;d\xi$ (47a)
$=\;\int_{0}^{1}\,\xi^{\text{B}_{0}+p}\,(1-\xi)^{\text{B}_{1}}\,e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi$ (47b)
$=\;e^{\text{B}_{2}}\,\Gamma(\text{B}_{0}+p+1)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (47c)
In Eq. (47a), we have effectively defined $\left<\xi^{p}\right$> to be $\langle{1}\,|\,{\xi^{p}}\rangle$}. In Eq. (47b), we have applied the definition of the scalar product, and in Eq. (47c), we have equated the integral over $\xi$ with the product of special functions found in Eq. (43), but with $\text{B}_{0}\rightarrow\text{B}_{0}+p$. The result is that the scalar product of any two smooth functions may be equated with a sum over confluent hypergeometric terms,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;$ $=\;e^{\text{B}_{2}}\,\sum_{j,k}\;\text{a}_{j}\,\text{b}_{k}\,\Gamma(\text{B}_{0}+j+k+1)\,$ (48)
$\hskip 56.9055pt\times\,U(1+\text{B}_{0}+j+k,-\text{B}_{1},-\text{B}_{2})\;.$
In practice, it is useful to precompute monomial moments using Eq. (47), and then compute scalar products using Eq. (46). The monomial moments are defined up to an overall constant. Henceforth, we chose to normalize all $\left<\xi^{p}\right$> such that $\left<\xi^{1}\right>=1$,
$\left<\xi^{p}\right>\;\leftarrow\;\left<\xi^{p}\right>/\left<\xi^{1}\right$>\;. (49)
VI Limitations of direct integration and analytic continuation
Direct integration, analytic continuation, and their application to monomial moments and confluent Heun polynomials are all subject to various caveats and limitations. Here, these facets are reviewed. At times, technical details are omitted for brevity, but particular attention is given to practical workarounds. While analytic continuation and direct integration are equivalent when their respective caveats are satisfied, they are otherwise found to be complementary. Ultimately, analytic continuation is determined to be easier to use in practice.
VI.1 Direct Integration: disallowed BH spins, and the zero frequency limit
The effective weight function introduced in Sec. V.1 was
$\text{W}_{\text{eff.}}(z)\;$ $=\;e^{\frac{\text{B}_{2}\text{x}_{2}}{z^{\text{x}_{0}}-1}}\;z^{\text{x}_{0}(\text{B}_{0}+1)-1}\;\left(1-z^{\text{x}_{0}}\right)^{\text{B}_{1}}\;$ (50)
$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
In Sec. V.1, a direct integration approach was introduced to evaluate scalar product using $\text{W}_{\text{eff.}}(z)$. This approach has two limiting features: (i) for a given value of $\tilde{\omega}$, there are values of the BH spin, $a$, such that an integration path cannot be constructed, and (ii) when $\tilde{\omega}=0$, the weight function may diverge at $\xi=1$.
For any value of $\tilde{\omega}=\omega-i/\tau$, there may exist values of the BH spin,
$a=a_{*}\;,$ (51)
such that an integration path cannot be constructed as described in Sec. V.1. This limitation results from the requirement that $\text{Re}(\text{x}_{0})>0$.
In Sec. V.1, the coordinate parameter $\text{x}_{0}$ was introduced so that $\text{W}_{\text{eff.}}(z)$ is zero at $z=0$. The value found for $\text{x}_{0}$ to have that effect was,
$\text{x}_{0}\;$ $=\;\text{K}_{0}\,\frac{\text{sgn}(\mathrm{Re}(\text{B}_{0})+1)}{\mathrm{Re}(\text{B}_{0})+1}$ (52)
$\quad\;+\;i\,\frac{-1\,+\,\text{K}_{0}\,[\text{sgn}(\;\mathrm{Re}(\text{B}_{0})+1\;)-1]}{\mathrm{Im}(\text{B}_{0})}$
where $\text{sgn}(x)$ is defined by $x=\text{sgn}(x)\,|x|$, and
$\text{B}_{0}\;=$ $\;-\left(\frac{2M(\delta+M)}{\tau\delta}+s\right)$ (53)
$\hskip 14.22636pt\;+\;i\left(\frac{am-2M^{2}\omega}{\delta}-2M\omega\right)\;.$
and $\text{K}_{0}>0$. As seen in Eq. (52), the real part of $\text{x}_{0}$ is generally regular, since $\text{sgn}(\mathrm{Re}(\text{B}_{0})+1)$ is defined to be zero when $\text{Re}(\text{B}_{0})+1$ is zero. However, the imaginary part of $\text{x}_{0}$ may diverge if $\text{Im}(\text{B}_{0})=0$. If we had constructed $\text{x}_{0}$ differently, e.g. by allowing $\text{Re}(\text{x}_{0})$ to depend on both $\text{Re}(\text{B}_{0})$ and $\text{Im}(\text{B}_{0})$, then the requirement that $\text{Re}(\text{x}_{0})>0$ would still have resulted in a $1/\text{Im}(\text{B}_{0})$ divergence.
Taking $\omega$ as known, there may exist values of the BH spin, namely $a_{*}$, such that
$\text{Im}{(\text{B}_{0})}|_{a=a_{*}}\;$ = $=\;0$ (54a)
Applying the definition of $\text{B}_{0}$ (see Eq. 53), and recalling that $\delta=\sqrt{M^{2}-a^{2}}$ allows Eq. (54a) to be solved for $a_{*}$,
$a_{*}\;$ $=\;\frac{4mM^{2}w}{m^{2}+4M^{2}w^{2}}\;.$ (54b)
For physical applications, it happens that for any QNM with $(\ell,m,n)$, having frequency
$\tilde{\omega}(a)\;=\;\omega(a)-i/\tau(a)\;,$ (55)
there is a discrete set BH spins where $\omega$ is such that $a=a_{*}$. For a given QNM $\ell$, $m$ and $n$, this set may be defined by first evaluating $\omega(a)$ using e.g. positive.physics.qnmobjLondon et al. (2020), and then searching for values of $a$ such that
$|a-a_{*}(\tilde{\omega}{(a)})|\;=\;0\;.$ (56)
For $(\ell,m,n)=(2,2,0)$, there is only one such BH spin,
$a_{*}\;\approx\;0.996632\;.$ (57)
It is instructive to contemplate the behavior of $\text{x}_{0}$ when $a$ is above this critical value.
Towards extremal BH spin, $a\rightarrow 1$, $\text{x}_{0}$ typically remains regular, meaning that an integration path, $\xi(z)$, may still be constructed. This is a result of how $\text{x}_{0}$ depends on $\text{B}_{0}$. Since $\lim_{a\rightarrow 1}|\text{Re}(\text{B}_{0})|\rightarrow\infty$, $\text{Re}(\text{x}_{0})$ tends to zero in that same limit. Similarly, it may be deduced that $\text{Im}(\text{x}_{0})$ tends to zero in that limit. Recall from Eq. (34) that the integration path, $\xi(z)$, is not defined when $\text{x}_{0}=0$. Despite this, the reasoning above implies that $\xi(z)$ may be constructed in the near extremal limit, within the limits of numerical precision; i.e. when $\text{x}_{0}$ is small, but strictly non-zero.
Lastly, a second coordinate parameter, $\text{x}_{2}$, was introduced so that exponential divergence of the weight function was avoided (see Sec. V.1). The full expression for $\text{x}_{2}$ was
$\text{x}_{2}\;$ $=\;\frac{\text{K}_{2}\,\text{x}_{0}}{\text{B}_{2}}\;$ (58a)
$=\;\frac{\text{K}_{2}\,\text{x}_{0}}{4i\delta\tilde{\omega}}\;$. (58b)
In Eq. (58b), we have used the fact that $\text{B}_{2}=4i\delta\tilde{\omega}$. Since $\text{x}_{2}$ is proportional to $\tilde{\omega}^{-1}$, it generally diverges in the zero frequency limit. However, since only $\text{x}_{2}\,\text{B}_{2}\,=\,\text{K}_{2}\,\text{x}_{0}$ appears in the weight function (see e.g. Eq. 50), $\text{W}_{\text{eff.}}(z)$ may still be evaluated.
Further, in the zero frequency limit there is no exponential divergence of the weight function at $\xi=1$, and so there is no formal need for $\text{x}_{2}$ at all in that limit. Seeing as
$\text{W}_{\text{eff.}}(z)|_{\tilde{\omega}=0}\;\propto\;\left(1-z^{\text{x}_{0}}\right)^{\text{B}_{1}}\;,$ (59)
$\text{W}_{\text{eff.}}$ will have a pole at $z=1$ if $\text{Re}(\text{B}_{1})<0$. This is of no concern when $\tilde{\omega}\neq 0$, since $\text{x}_{2}$ is designed to enforce exponential damping at $z=1$; otherwise, when $\tilde{\omega}=0$, it is possible to construct a different coordinate, $z^{\prime}$, such that this pole is avoided. This work-around may only be viable when $\tilde{\omega}=0$; otherwise, it is found that when $\tilde{\omega}\neq 0$, the required coordinate change necessarily introduces oscillations in $\xi(z^{\prime})$ that grow exponentially as $z$ increases. Thus, when $\tilde{\omega}=0$, one can only use coordinate freedom to either avoid the essential singularity at $\xi(z)=1$ (i.e. that resulting from the exponential), or the pole at $\xi(z)=1$, but not both.
In this, direct integration’s underlying limitation is one of mathematical perspective: it happens to be more robust to use analytic continuation via integration over closed contours, rather than singularity avoiding integration over finite paths as has been discussed above. For example, Ref. Ronveaux and Arscott (1995) details how single and double Pochhammer contours may be used in similar mathematical settings. This approach makes use of standard theorems in complex analysis, and is ultimately equivalent to analytic continuation as outlined in Sec. V.2.
VI.2 Analytic continuation and limitations of the Gamma function
It was found in Sec. V.2 that scalar products may be evaluated analytically using monomial moments, $\left<\xi^{p}\right$>,
$\left<\xi^{p}\right$>\; $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (60a)
The qualitative behavior of the monomial moments can be described by
$\left<\xi^{p}\right$>\; $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}\;,$ (60b)
where the leading term in the power series definition for $U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})$ has been used to approximate it with ${\Gamma(1+\text{B}_{1})}/{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}$ DLMF (a). Since $\text{B}_{2}=4i\delta\tilde{\omega}{}$, Eq. (60b) is exact in the zero frequency limit, $\tilde{\omega}=0$, or in the extremal spin limit, $\delta=\sqrt{M^{2}-a^{2}}=0$.
For either Eq. (60a) or Eq. (60b), $\left<\xi^{p}\right$> will be regular when instances of the Gamma function are regular. Since $\Gamma(z)$ is well known to have simple poles when $z$ is a negative integer, it follows that $\left<\xi^{p}\right$> can diverge when,
$1+\text{B}_{0}+p$ $\in\{\;z\in\mathbb{Z}\;|\;z<0\;\}\;,$ (61a)
and
$1+\text{B}_{1}$ $\in\{\;z\in\mathbb{Z}\;|\;z<0\;\}\;.$ (61b)
In Eq. (61), $\{z\in\mathbb{Z}\,|\,z<0\}$ denotes the set of all negative integers. The practical implications of Eq. (61) may be investigated by rewriting $\text{B}_{0}$ and $\text{B}_{1}$ in terms of physical parameters,
$p-s+1-\frac{2M(\delta+M)}{\tau\,\delta}\quad\quad\quad\quad\;\;$ (62a)
$+\;\frac{i(am-2M\omega(\delta+M))}{\delta}$ $\;=\;-z_{0}\;,$
$2s+1-\frac{4M}{\tau}-4iM\omega$ $\;=\;-z_{1}\;.$ (62b)
In Eq. (62), $z_{0}$ and $z_{1}$ are natural numbers, and we have used Eq. (8), Eqs. (16) and (24) to expand in terms of basic physical quantities. Equation (62a) is equivalent to Eq. (61a), and Eq. (62b) is equivalent to Eq. (61b).
Inspection of Eqs. (62a) and (62b) reveals that they can only hold when $M\tilde{\omega}{}$ is purely imaginary (i.e. $\omega=0$). This limits the physical relevance of these equations to the total transmission QNMs Cook and Zalutskiy (2014). Furthermore, Eq. (62a) can only be true if $a=0$, thus it is only relevant to the algebraically special modes of Schwarzschild Cook and Zalutskiy (2014); Berti et al. (2003).
Both Eqs. (62a) and (62b) have non-trivial implications for evaluation of the scalar product with analytic continuation. If Eq. (62a) holds, then there exist some monomials, $\xi^{p}$, such that the respective monomial moments are infinite when defined by Eq. (60a). If Eq. (62b) holds, then no monomial moments are finite according to Eq. (60a).
These limitations are, again, a result of mathematical perspective: evaluation of the scalar product with Gamma and confluent hypergeometric functions may fail, depending on $\text{Re}(\text{B}_{0})$ or $\text{Re}(\text{B}_{1})$; however, we saw in the previous section that the viability of direct integration only depends on $\text{Im}(\text{B}_{0})$. Thus the two perspectives have complementary limitations.
VI.3 Direct Integration vs. Analytic continuation
In practice, analytic continuation has the following relative attributes that may make it preferable in most circumstances. Since analytic continuation does not require the determination of optimal coordinate parameters (see Sec. V.1), its implementation requires fewer steps. In addition, analytic continuation requires monomial moments to compute scalar products, ans do does not require the storage of arrays (discretizations of $\xi(z)$) of length $10^{5}$ or greater (see Sec. V.1). Instead, monomial moments with degrees up to twice that of the largest polynomial order of interest may be precomputed, ideally using arbitrary precision arithmetic mpmath development team (2023), and then used for all scalar products (Sec. V.2). Thus, while direct integration may serve as a valuable alternative, particularly in cases where the Gamma function cannot be evaluated, analytic continuation is expected to be structurally simpler and computationally more efficient.
All forthcoming numerical results will make use of analytic continuation with $64$ digits of precision via the public mpmath package mpmath development team (2023).
Refer to caption Refer to caption Refer to caption
Figure 3: Example distributions of absolute monomial moments, $|\langle\xi^{k}\rangle|$ v.s. monomial order, $k$, for the $(\ell,m)=(2,2)$ QNM at BH dimensionless spin of $a/M=0.7$. Left, central and right panels show results for $n$ of 0, 22 and $6$, respectively. The corresponding QNM frequencies are $M\tilde{\omega}=0.5326-0.0808i$, $M\tilde{\omega}=0.4999-0.4123i$, and $M\tilde{\omega}=0.4239-1.0954i$ respectively. At focus is the effect of negating the spin weight, $s$, for increasing overtone label. For all cases shown, $|\langle\xi\rangle|=1$ has been enforced. Closed circles show points for $s=+2$, and open circles show points for s=2s=-2.
VII Pedagogical Applications
Previous sections provide a foundation for applying the scalar product, Eq. (63), to physical scenarios in which the series expansion for radial functions is either known (analytic continuation), or unknown (direct integration). Direct integration is applicable in both scenarios.
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (63)
Before leaping into applications to Teukolsky’s radial equation, there are good reasons to first look closely at two pedagogical cases.
In this section, we will first apply the scalar product to the monomial moments for different QNM values of the frequency $\tilde{\omega}$, paying particular attention to the effect of spin weight, $s$. For a broad range of scenarios, monomial moments with $s=2$ decrease monotonically with monomial order, while this is less so the case for s=2s=-2. We will then apply the scalar product to the confluent Heun polynomials.
Previously, Cook and Zalutskiy considered confluent Heun polynomials for BHs in Ref. Cook and Zalutskiy (2014); Fiziev (2009a). There, the polynomials were defined such that a special choice $\tilde{\omega}{}$ resulted in $\mathcal{L}_{\xi}$ (Eq. 15) having polynomial eigenfunctions. These are the well known algebraically special and total transmission QNMs Cook and Zalutskiy (2014); Andersson (1994); Dias and Reall (2013); Berti et al. (2003); Whiting (1989); Fiziev (2009a).
Here, we take a different perspective: We will a priori define $\tilde{\omega}{}$ to be any QNM frequency  [^15] In the language of Ref. Ronveaux and Arscott (1995), we choose to focus on Class I of confluent Heun polynomials.. This requires that, in effect, we redefine the potential of $\mathcal{L}_{\xi}$ such that its eigenfunctions are polynomials. The consequences of this choice enable confluent Heun polynomials to be defined for arbitrary frequencies; in turn, this will have a significant impact on forthcoming results and discussion.
We will use the scalar product to derive the orthogonality between confluent Heun polynomials of like order. Lastly, we will use the scalar product to derive a relationship between the confluent Heun polynomials’ eigenvalues, and a specific average of $\xi$. This relationship provides some insight into the structure of the eigenvalues.
VII.1 Monomial moments
The scalar-product of any two smooth functions may be evaluated using a linear combination of monomial moments,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\sum_{j,k}\;\text{a}_{j}\,\text{b}_{k}\,\left<\xi^{j+k}\right$>\;. (64)
For many common scalar products, monomial moments $\left<\xi^{p}\right$> decrease monotonically with $p$. This is easily demonstrated by considering a hypothetical constant weight function:
$\text{W}(\xi)_{\text{hypothetical}}\;$ = $=\;1\;,$ (65a)
${\langle{\text{a}}\,|\,{\text{b}}\rangle}_{\text{hypothetical}}\;$ $=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)d\xi\;,$ (65b)
$\left<\xi^{p}\right>_{\text{hypothetical}}\;$ $=\;(p+1)^{-1}\sim p^{-1}\;.$ (65c)
In Eq. (65a), a weight function of $1$ results in a scalar product equivalent to that used for the shifted Legendre polynomials, Eq. (65b). In turn, this results in monomial moments that are asymptotically equivalent to one over the degree of the monomial, Eq. (65c). In practice, this means that if the sequences comprised of $\text{a}_{j}$ and $\text{b}_{k}$ are convergent, then so too is $\text{a}_{j}\,\text{b}_{k}\,\left<\xi^{j+k}\right$>. We will now examine this argument in the context of the QNM’s scalar product. Our goal is to gain a basic and qualitative understanding of how monomial moments depend on two physical parameters: $s$ and $M\tilde{\omega}$.
We previously observed that monomial moments may be computed via analytic continuation of the confluent hypergeometric function, $U$, and the gamma function, $\Gamma$,
$\left<\xi^{p}\right$>\; $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;$ (66)
The confluent hypergeometric function, $U$, is defined as a series expansion in its third input DLMF (a). Thus some understanding of $\left<\xi^{p}\right$> may be found by examining $U$ expanded in powers of $\text{B}_{2}$. Examination of the zeroth order term happens to be sufficient for qualitative understanding of how $\left<\xi^{p}\right$> behaves when $p$ is large. For precise and predictive understanding, one would have to extend the forthcoming analysis to higher orders. This will not be done here so that the following discussion remains fairly simple.
To zero’th order in its third input, $U$ takes the approximate form,
$U(\text{x},\text{y},\text{z})\;\approx\;\frac{\Gamma(1-\text{y})}{\Gamma(\text{x}-\text{y}+1)}+\mathcal{O}(z)\;.$ (67)
When applied to Eq. (66), this yields
$\left<\xi^{p}\right$>\; $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}$ (68a)
$\sim\;p^{-1-\text{B}_{1}}\;\Gamma(1+\text{B}_{1})\;.$ (68b)
To go from Eq. (68a) to Eq. (68b), we have used the fact that, as $p\rightarrow\infty$, $\Gamma(p+\alpha)\,\sim\,p^{\alpha}\,\Gamma(p)$ Abramowitz and Stegun (1964). The physics of the situation is revealed by using Eqs. (24) and (16) to rewrite $\text{B}_{1}$ in terms of $M$, $\tilde{\omega}$ and $s$,
$\left<\xi^{p}\right$>\; $\sim\;p^{-1-2s+4iM\tilde{\omega}}\;$ (69a)
$\sim\;p^{-1-2s+4M/\tau}e^{4iM\omega\,\ln(p)}\;.$ (69b)
[…]
The presence of $e^{4iM\omega\,\ln(p)}$ in Eq. (69b) means that, in general, the real and imaginary parts of $\left<\xi^{p}\right$> will trace out a logarithmic spiral.
The presence of $p^{-1-2s+4M/\tau}$ within Eq. (69b) allows us to discern three additional features:
  1. (i)
    There exist some combination of $s$ and $M/\tau$ such that monomial moments either decrease, asymptote to a constant, or increase as $p$ tends towards infinity.
  2. (ii)
    Since Eq. (69b) depends on $p^{-2s}$, monomial moments for $s=+2$ will be much more inclined to decrease as $p\rightarrow\infty$:
    $\left<\xi^{p}\right>|_{s=+2}\;\sim\;\ p^{-8}\,\left<\xi^{p}\right>|_{s=-2}$ (70)
    In Eq. (70), we have used the fact that if $\tilde{\omega}{}$ is a QNM frequency, then $\tilde{\omega}|_{s=+2}=\tilde{\omega}|_{s=-2}$.
  3. (iii)
    Lastly, the presence of $4M/|\tau|$ in Eq. (69b) means that, as $\tau$ decreases, the rate of change of $\left<\xi^{p}\right$> will typically increase. This is perhaps most relevant for the QNM overtones, for which $|\tau|$ decreases as the overtone label, $n$, increases Zimmerman and Mark (2016); Berti et al. (2003).
These three qualitative features are visible in Fig. 3:
  1. (i)
[…]
  2. (ii)
    In each panel, we see that monomial moments for spin weight $s=+2$ are significantly depressed relative to those for s=2s=-2.
  3. (iii)
    The scale of each panel is identical so that they effect of increasing overtone label (i.e. decreasing $\tau$) is easily visible for both spin weights.
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: The first four confluent Heun polynomials for $p=k$, spin weight s=2s=-2, black hole dimensionless spin $a/M=0.86$, and $\tilde{\omega}$ defined by the $(\ell,m,n)=(3,3,1)$ QNM. The corresponding QNM frequency is $M\tilde{\omega}=0.9840-0.2140i$. Polynomial orders $p\in\{0,1,2,3\}$ are shown as a function of the physically valued fractional radius $\xi$. The top left and right panels shows the real and imaginary parts of each polynomial, respectively. Bottom left and right panels show the corresponding polynomial amplitudes and phases, respectively.
[…]
Since the weight function, $\text{W}(\xi)$, allows the computation of arbitrary monomial moments, the monomials are a basis for the scalar product space defined by Eq. (64Adkins and Weintraub (2012); Gohberg et al. (2006). Thus, it is fair to wonder whether there exists a related basis that is more suitable for numerical approximation? This is the basic question that links the study of monomial moments to applied polynomial theory Chihara (2011). A complete investigation of this question is beyond the current scope. Nevertheless, a first step towards an answer may be found by applying the scalar product to confluent Heun polynomials Whiting (1989); Cook and Zalutskiy (2014); Fiziev (2009b); DLMF (b); Magnus et al. (2021); Fiziev (2009a).
VII.2 Confluent Heun polynomials
The differential operator that results from applying QNM boundary conditions to Teukolsky’s radial equation was
$\mathcal{L}{{}_{\xi}}\;$ $=\;(\text{C}_{0}+\text{C}_{1}(1-\xi))$ (71a)
$+\;\left(\text{C}_{2}+\text{C}_{3}(1-\xi)+\text{C}_{4}(1-\xi)^{2}\right)\partial_{\xi}$ (71b)
$+\;\xi(\xi-1)^{2}\partial_{\xi}^{2}\;.$ (71c)
Since the operator given in Eq. (71) is of the confluent Heun type, its eigen-solutions are the aptly named confluent Heun functions Ronveaux and Arscott (1995); Magnus et al. (2021).
The subset of these with terminating series solutions are, unsurprisingly, named confluent Heun polynomials DLMF (a); Cook and Zalutskiy (2014); Fiziev (2009b); Ronveaux and Arscott (1995).
A confluent Heun polynomial of order $p$ is typically considered to result from special values of $\text{C}_{0}$ and $\text{C}_{1}$ Cook and Zalutskiy (2014); Fiziev (2009b). In the context of Teukolsky’s radial equation, these special values have been interpreted as a constraint on $\tilde{\omega}$. This is the case for the algebraically special modes of Schwarzschild and the total transmission modes of Kerr Cook and Zalutskiy (2014); Berti et al. (2003); Fiziev (2009a). Here, a different perspective on the confluent Heun polynomial conditions for BHs will be used. This perspective is based on two observations:
  1. (i)
    Classical polynomials are typically defined to be eigenfunctions of a purely differential operator (i.e. one with no $\partial_{\xi}^{0}$ term). For the confluent Heun operator, Eq. (71), this purely differential part is denoted by Eqs. (71b) and (71c), and repeated below in Eq. (72),
    $\mathcal{D}{{}_{\xi}}\;$ $=\;\left(\text{C}_{2}+\text{C}_{3}(1-\xi)+\text{C}_{4}(1-\xi)^{2}\right)\partial_{\xi}$ (72)
    $\hskip 14.22636pt+\;\xi(\xi-1)^{2}\partial_{\xi}^{2}\;.$
  2. (ii)
    Any attempt to construct order $p$ polynomials that are eigenfunctions of $\mathcal{D}_{\xi}$ requires that the “eigenvalue” be a linear function of the domain variable,
    $\mathcal{D}{{}_{\xi}}\;\mathbf{y}_{pk}(\xi)\;$ $=\;\sigma_{pk}(\xi)\;\mathbf{y}_{pk}(\xi)\;,$ (73a)
    $\sigma_{pk}(\xi)\;$ $=\;\lambda_{pk}\;+\;\mu_{p}\,\xi\;.$ (73b)
    Plainly, $\sigma_{pk}(\xi)$ is not an eigenvalue in the traditional sense. Instead, it is an instance of a two parameter eigenvalue problem Ronveaux and Arscott (1995); Atkinson and Mingarelli (2011). We will see that $\mu_{p}$ is uniquely determined by the order of the polynomial, meaning that it need not be conceptually separated from the traditional eigenvalue, $\lambda_{pk}$.
In what follows, we will apply these ideas to construct confluent Heun polynomials that are valid for any value of $\tilde{\omega}$ (i.e. any QNM frequency). We will see that the operator, $\mathcal{D}{{}_{\xi}}$, is self-adjoint under the same scalar product that we have discussed previously. We will then use the self-adjointness of this operator to prove and numerically demonstrate the orthogonality of confluent Heun polynomials for fixed polynomial order. Lastly, we will use self-adjointness to derive an expression for the confluent Heun polynomials’ eigenvalues in terms of ratios of scalar products.
VII.2.1 Construction
Let $\mathbf{y}_{pk}(\xi)$ be a confluent Heun polynomial of order $p$,
$\mathbf{y}_{pk}(\xi)\;=\;\sum_{q=0}^{p}\,\text{a}_{pkq}\,\xi^{q}\;.$ (74)
We will soon see that the label, $k$, is useful because there are will generally be $p+1$ order $p$ polynomials.
The polynomial coefficients, $\text{a}_{pkq}$, are determined by applying the form above for $\mathbf{y}_{pk}(\xi)$ to the eigen-relationship given by Eq. (73a). One then asserts that the coefficients of each power in $\xi$ must be equal on both sides of Eq. (73a). The result is a three term recursion relationship or, equivalently, a discrete eigenvalue relationship. To simplify our notation, we will henceforth drop two indices, $p$ and $k$, from $\text{a}_{pkq}$,
$\text{a}_{q}\;\leftarrow\;\text{a}_{pkq}\;.$ (75)
Thus the discrete eigenvalue relationship takes the schematic form,
$\text{a}_{q-1}\,\alpha_{q}\;+\;\text{a}_{q}\,\beta_{q}\;+\;\text{a}_{q+1}\,\gamma_{q}\;=\;\lambda_{pk}\,\text{a}_{q}\;,$ (76)
where
$\alpha_{q}\;$ $=\;-(q-1)(\text{C}_{4}+q-2)-\mu_{p}$ (77a)
$\beta_{q}\;$ $=\;q\,(\text{C}_{3}+2(\text{C}_{4}+q-1))$ (77b)
$\gamma_{q}\;$ $=\;-(q+1)(\text{C}_{2}+\text{C}_{3}+\text{C}_{4}+q)\;.$ (77c)
An initial boundary condition,
$\text{a}_{0}\,\beta_{0}\;+\;\text{a}_{1}\,\gamma_{0}\;=\;\lambda_{pk}\,\text{a}_{0}\;,$ (78)
is imposed for consistency with Eq. (74). In Eq. (78), $\text{a}_{0}=1$ is a common initial normalization choice.
A necessary but insufficient condition for the three-term recursion to terminate at the $p^{\text{th}}$ term, is simply that $\alpha_{p+1}=0$. This is the so-called $\Delta_{p+1}=0$ condition DLMF (b); Cook and Zalutskiy (2014). In our notation, this takes the form,
$\alpha_{p+1}\;$ $=\;p\,(\,p+\text{C}_{4}-1)-\mu_{p}\;=\;0\;.$ (79)
In turn, this can be interpreted as a constraint that allows the determination of $\mu_{p}$,
$\mu_{p}\;$ $=\;p\,(\,p+\text{C}_{4}-1\,)\;.$ (80)
Thus $\mu_{p}$ must be non-zero for the existence of non-trivial polynomial solutions to Eq. (73a).
Furthermore, if $\mu_{p}$ is given by Eq. (80), then the discrete eigenvalue relationship given by Eq. (76) defines a finite dimensional matrix eigenvalue problem [^17] Note that, along with the series’ boundary conditions (Eqs. 78-79), it is the finite nature of the system of equations that justifies the eigenproblem interpretation. In contrast, non-polynomial solutions to the radial problem will not satisfy Eq. (79), meaning that e.g. a continued fraction based method, and a related boundary condition at the infinite end of the series are required Leaver (1986, 1985).  Ronveaux and Arscott (1995); DLMF (b); Cook and Zalutskiy (2014). The matrix of interest, $\hat{\mathbf{A}}_{p}$, is asymmetric and tridiagonal,
$\hat{\mathbf{A}}_{p}\;=\;\left(\begin{array}[]{ccccc}\beta_{0}&\gamma_{0}&0&\cdots&0\\ \alpha_{1}&\beta_{1}&\gamma_{1}&\ddots&0\\ 0&\alpha_{2}&\beta_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\gamma_{p-1}\\ 0&0&\cdots&\alpha_{p}&\beta_{p}\end{array}\right)\;.$ (81)
By construction, the eigenvectors of $\hat{\mathbf{A}}_{p}$ are the polynomial coefficients, and the eigenvalues of $\hat{\mathbf{A}}_{p}$ are the polynomial eigenvalues,
$\{\text{a}_{pkq}\}_{q=0}^{p}\;$ $=\;{k}^{\text{th}}\text{ Eigenvector of }\hat{\mathbf{A}}_{p}\;,$ (82a)
$\lambda_{pk}\;$ $=\;{k}^{\text{th}}\text{ Eigenvalue of }\hat{\mathbf{A}}_{p}\;.$ (82b)
Since $\hat{\mathbf{A}}_{p}$ is a $(p+1)\times(p+1)$ matrix, there will be $p+1$ eigenvalues, $\lambda_{pk}$. If the eigenvalues are unique, then we may impose an order on $\mathbf{y}_{pk}$, whereby
$k\in\{0,1,2,...,p\}\;.$ (83)
In general, the $p+1$ values of $\lambda_{pk}$ will correspond to the roots of $\hat{\mathbf{A}}_{p}$’s characteristic polynomial. Lacking a deeper understanding how the roots of all such polynomials are distributed, the following ordering will henceforth be used
$\text{Im}\,\lambda_{p0}\,<\,\text{Im}\,\lambda_{p1}\,<\,\text{Im}\,\lambda_{p2}\,<\,...\,<\,\text{Im}\,\lambda_{pp}\;.$ (84)
In this sense, for a given order $p$, there is a multiplex comprised of $p+1$ solutions, where each solution is labeled by the multiplex index, $k$.
VII.2.2 Orthogonality between polynomials of like order
If each $\lambda_{pk}$ is unique, then each corresponding polynomial, $\mathbf{y}_{pk}$, will be linearly independent Axler (2015). Since $\mathcal{L}_{\xi}=\mathcal{D}_{\xi}+(\text{C}_{0}+(1-\xi)\text{C}_{1})$ is self-adjoint with respect to the scalar product, Eq. (64), $\mathcal{D}_{\xi}$ is also self-adjoint,
$\mathcal{D}{{}_{\xi}}\;$ $=\;{\mathcal{D}{{}_{\xi}}}^{\dagger}\;$. (85)
Here, we will use the self-adjointness of $\mathcal{D}_{\xi}$ to show that when $\lambda_{pk}$ are unique, $\mathbf{y}_{pk}$ must be orthogonal. To that end we will use a standard argument that is based on the behavior of $\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathcal{D}_{\xi}\mathbf{y}_{pk}}\rangle$} when $(p^{\prime},k^{\prime})$ swaps places with $(p,k)$. Along the way, we will encounter a twist that results from the fact that $\mathbf{y}_{pk}$ satisfy a non-standard eigenvalue equation, Eq. (73a).
Let us begin by analyzing the matrix whose elements are $\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathcal{D}_{\xi}\mathbf{y}_{pk}}\rangle$}. Since $\mathcal{D}_{\xi}$ is self-adjoint with respect to the scalar product (Eq. 26), we have that
${\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathcal{D}{{}_{\xi}}\,\mathbf{y}_{pk}}\rangle}\;$ $=\;{\langle{\mathcal{D}{{}_{\xi}}\,\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;.$ (86)
Applying the eigenvalue relationship,
$\mathcal{D}{{}_{\xi}}\;\mathbf{y}_{pk}(\xi)\;$ $=\;(\lambda_{pk}\;+\;\mu_{p}\,\xi)\;\mathbf{y}_{pk}(\xi)$ (87)
to both sides of Eq. (86) yields
$\lambda_{pk}$ ${\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;+\;\mu_{p}\,\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}|\xi\;|{\mathbf{y}_{pk}}\rangle$ (88)
$=\;\lambda_{p^{\prime}k^{\prime}}{\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;+\;\mu_{p^{\prime}}\,\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}|\xi\;|{\mathbf{y}_{pk}}\rangle\;.$
A useful rearrangement of Eq. (88) is given by Eq. (89),
$(\lambda_{pk}-\lambda_{p^{\prime}k^{\prime}})\,$ ${\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;=\;(\mu_{p^{\prime}}-\mu_{p})\,\langle{\mathbf{y}_{p^{\prime}k^{\prime}}}|\xi\;|{\mathbf{y}_{pk}}\rangle\;.$ (89)
The aforementioned twist is that, unlike in single parameter eigenvalue problems, the right-hand-side of Eq. (89) is non-zero. Nevertheless, the standard orthogonality argument is recovered by restricting Eq. (89) to polynomials of like order,
$p\;$ $=\;p^{\prime}\;.$ (90)
The resulting version of Eq. (89) is,
$(\lambda_{pk}-\lambda_{pk^{\prime}})\,{\langle{\mathbf{y}_{pk^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;$ = $=\;0\;.$ (91)
If $\lambda_{pk}$ are distinct for different $k$, then Eq. (91) can only hold if
${\langle{\mathbf{y}_{pk^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;$ $=\;\eta_{pk}^{2}\,\delta_{k^{\prime}k}\;.$ (92)
In Eq. (92), $\eta_{pk}$ is an inverse normalization constant. Henceforth, we will work with the normalized confluent Heun polynomials,
${y}_{pk}\;$ $=\;\eta_{pk}^{-1}\;\mathbf{y}_{pk}\;,$ (93)
where,
${\langle{{y}_{pk^{\prime}}}\,|\,{{y}_{pk}}\rangle}\;$ $=\;\delta_{k^{\prime}k}\;.$ (94)
Refer to caption Refer to caption Refer to caption
Figure 5: Distribution of roots for confluent Heun polynomials, $y_{pk}(\xi)$, for three cases, all with s=2s=-2, $a/M=0.7$, and $\ell=m=2$. The three cases differ in polynomial order, $p$, and the frequency, $\tilde{\omega}$, used: (left) $y_{33}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.5326-0.0808i$, was used, (center) $y_{77}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.4713-0.5843i$, was used, and (right) $y_{20,20}$, the $n=12$ QNM frequency, $M\tilde{\omega}=0.4155-2.5050i$, was used. Polynomial roots are marked by white open circles. Roots were calculated using numpy.roots Horn and Johnson (1990); Harris et al. (2020). Values of polynomial phase, $\arg(y_{pk})$, are shaded from blue to yellow, indicating values between $\pi$ and $\pi$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
VII.2.3 A scalar product expression for eigenvalues
For polynomials of unlike order, Eq. (89) may be recast as a relationship between polynomial eigenvalues. This is done by solving for $\lambda_{pk}$,
$\lambda_{pk}\;=\;\lambda_{p^{\prime}k^{\prime}}\,+\,(\mu_{p^{\prime}}-\mu_{p})\,\frac{\langle{{y}_{p^{\prime}k^{\prime}}}|\xi\;|{{y}_{pk}}\rangle}{{\langle{{y}_{p^{\prime}k^{\prime}}}\,|\,{{y}_{pk}}\rangle}}\;.$ (95)
Equation (95) applies to all $(p,k)$ and $(p^{\prime},k^{\prime})$, where $p\neq p^{\prime}$. An instructive special case is that of
$p^{\prime}=0\;.$ (96)
This corresponds to ${y}_{00}$, which must be constant. If ${y}_{00}$ is a constant, then for the eigenvalue relation (Eq. 87) to hold, it must be true that
$\lambda_{00}\;$ $=\;\mu_{0}\;=\;0\;.$ (97)
Applying Eq. (97) to Eq. (95) gives
$\lambda_{pk}\,\;$ $=\;\mu_{p}\,\frac{\langle{y_{00}}|\xi\;|{y_{pk}}\rangle}{{\langle{y_{00}}\,|\,{y_{pk}}\rangle}}\;.$ (98)
Recalling that $\mu_{p}=p\,(\,p+\text{C}_{4}-1\,)$, Eq. (98) is equivalent to
$\lambda_{pk}\,\;$ $=\;p\,(\,p+\text{C}_{4}-1\,)\,\frac{\langle{y_{00}}|\xi\;|{y_{pk}}\rangle}{{\langle{y_{00}}\,|\,{y_{pk}}\rangle}}\;.$ (99)
If $y_{pk}$ are known, then Eq. (99) allows the computation of the related eigenvalues via the scalar product. In most instances this will not be needed to define the eigenvalues; however, Eqs. (95-99) are presented to provide a way to check consistency between evaluations of the scalar product, and the underlying confluent Heun system.
VII.2.4 Numerical examples
Numerical computation of the confluent Heun polynomials may proceed by first computing $\text{C}_{2}$, $\text{C}_{3}$ and $\text{C}_{4}$. Then, for a chosen polynomial order, one would compute $\hat{\mathbf{A}}_{p}$ (Eq. 81), so that its eigenvalues and vectors may be determined. Since scalar products between polynomials is a weighted sum over monomial moments (Eq. 64), it is useful to precompute and store all of the monomial moments relevant to a given polynomial order. This allows all polynomial scalar products to be evaluated without explicit integration (Eq. 47). From this perspective, numerical evaluation of confluent Heun polynomials may be accomplished on-the-fly via the appropriate sum over monomials (Eq. 74). Here, this procedure is used to provide select numerical examples.
Figure 4 shows example confluent Heun polynomials of orders $p\in\{0,1,2,3\}$. Only the polynomials for which $p=k$ are shown; others have qualitatively similar behavior. As expected, the real and imaginary parts of each case behave as polynomials in $\xi$. However, since each polynomial has complex value coefficients, the polynomials’ amplitudes and phases are non-polynomial functions of $\xi$. Similarly, since the polynomials are complex valued, we should not expect the roots of their respective real and imaginary parts to be simultaneous. For example, the top left panel of Fig. 4 shows $\text{Re}\,y_{11}$ having a root near $\xi=0.5$, but $\text{Im}\,y_{11}$ having no root at all on the real line.
Figure 5 shows the distribution of polynomial roots for three cases. In focus is the distribution of roots in the complex plane as polynomial order and overtone label increase. As expected, roots are typically not real valued. For large $n$ and large polynomial orders, roots are found to form two families, one situated around the origin, and another dispersed across the upper-right quadrant. This is seen in the right panel mof Fig. 5. In all panels of Fig. 5, each root is seen along with its associated phase boundary. Any smooth curve, e.g. $\xi(z)$, may be defined to have a continuous phase by adding the appropriate integer factor of $2\pi$ at each phase boundary.
Figure 6 shows examples of orthogonality at fixed polynomial order,
$p\;=\;20\;.$ (100)
for three cases. At focus, is the effect of increasing QNM overtone label or, equivalent, decreasing the QNM decay time $\tau$, such that $1/\tau$ increases within $\tilde{\omega}=\omega-i/\tau$. The $p=20$ polynomials are considered to be somewhat extreme cases: for this and all lower polynomial orders, no appreciable numerical artifacts are encountered for the numerical and physical parameters used here.
The left columns of Fig. 6 show the absolute values of Gramian matrix elements, $\langle y_{pk^{\prime}}|y_{pk}\rangle$, and the right columns show absolute values of $\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle$. Scalar products have been evaluated using analytic continuation (Eq. 64) with $64$ significant figures of precision via the dps setting of mpmath mpmath development team (2023).
Orthogonality of the polynomials at fixed order is seen through the diagonal structure of each Gramian. The top left panel, having $n=0$, displays a slight deviation away from orthogonality (i.e. values near $10^{-40}$ rather than $10^{-50}$ and below) as $k$ nears it maximum value of $20$.
This effect has two causes, one technical and the other mathematical. The technical cause of this feature is simply the value of the precision parameter, dps=64=64, which means that numerical results near $10^{-64}$ are more prone to round-off error. The mathematical cause of this feature is that, like the legendre polynomials, the confluent Heun polynomials have normalization constants that generally decrease as polynomial order increases [^18] For the Legendre polynomials, the analog of $\eta_{pk}^{2}$ (see Eq. 92) is $(2p+1)^{-1}$.. Consequently the action of normalization effectively scales each polynomial by a very large number, exacerbating finite precision errors. At worst, square normalization constants might be $10^{-64}$ or smaller and would thereby be corrupted by finite precision error. In turn, this would correspond to normalization constants $10^{-32}$ or smaller, and the act of normalization would be to scale the non-normalized polynomials by a corrupted number of order $10^{32}$. This situation can be avoided by increasing numerical precision.
The confluent Heun polynomials, while orthogonal at fixed order, do not display a feature common among classical polynomials. In particular, they do not possess three-term-recursion relations between polynomials of different values of $k$ Chihara (2011); Arfken et al. (2013). This may be understood thusly: Since $\xi\times y_{pk}$ is a polynomial or order $p+1$, it is to be expected that it cannot be written in terms of only polynomials of order $p$. Consequently, the matrix whose elements are $\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle$ will in general have no non-zero values, and therefore will not be band-diagonal.
This feature of the confluent Heun polynomials is shown in the right column of Fig. 6. There, the structure of $|\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle|$ is seen to increase non-trivially with overtone label, $n$. This may be partially understood in the context of the polynomial eigenvalues.
It was previously noted that the small deviations from orthogonality are linked to small normalization constants, and that increasing structure in the distribution of $|\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle|$ are linked to the polynomial eigenvalues. Polynomial normalization constants and eigenvalues are shown, respectively, in the left and right panels of Fig. 7. There, the same three cases seen in Fig. 6 are considered.
At focus in the left panel of Fig. 7 is the distribution of normalization constants. While normalization constants are found to generally decrease with polynomial order (not shown in figure), they are also found to generally decrease as $k$ increases, and they are found to generally increase as the imaginary part $\tilde{\omega}$ increases.
At focus in the right panel of Fig. 7 is how the distribution of eigenvalues changes as the imaginary part of $\tilde{\omega}$ becomes more negative (i.e. as $n$ increases). For $n=0$, either $\text{Re}\,\tilde{\omega}$ or $\text{Im}\,\tilde{\omega}$ may be used to define equivalent orders (i.e. maps between polynomials and the index $k$). This is not the case for $n\in\{6,12\}$, where branches appear in the eigenvalue distribution. In these cases, the eigenvalues are still unique, so a well defined order can still be constructed.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Gramian matrices for three QNM overtone labels: $n=0$ (top row), $n=6$ (middle row), and $n=12$ (bottom row). Considered here are the order $p=20$ normalized confluent Heun polynomials with $s=2$, $a/M=0.70$ and QNM frequency labels $(\ell,m)=(2,2)$. The QNM frequencies for top middle and bottom rows are $M\tilde{\omega}=0.5326-0.0808i$, $M\tilde{\omega}=0.4239-1.0954i$ and $M\tilde{\omega}=0.4098-3.1270i$, respectively. The left column shows the absolute values of $\langle y_{pk^{\prime}}|y_{pk}\rangle$; there, note that the color bar is log scaled between $10^{-50}$ and $10^{0}$. The right column shows absolute values of $\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle$; there, the color bar is logs scaled between $10^{-12}$ and $10^{0}$ to accentuate graph structure.
Refer to caption Refer to caption
Figure 7: Distribution of normalization constants, $\eta_{pk}$, and eigenvalues, $\lambda_{pk}=\mathrm{Re}\lambda_{pk}+i\,\mathrm{Im}\lambda_{pk}$, for three different QNM overtone labels: $n\in\{0,6,12\}$. For all values of $n$, the spin weight is $s=2$, the black hole dimensionless spin is $a/M=0.70$, and the QNM $(\ell,m)$ are $(2,2)$. The left panel shows normalization constants, and the right panel shows eigenvalues. Black dots mark $n=0$, orange circles mark $n=6$, and magenta squares mark $n=12$.
VIII Implications for analytic solutions to Teukolsky’s radial equation
We previously set out to apply two concepts to Teukolsky’s radial problem. When stated in the context of the radial equation, these two concepts are:
  • (i)
    That a scalar product may be developed from the structure of the radial equation (i.e. (Q1)).
  • (ii)
    That special functions may be developed from the structure of the radial equation, and that these functions may be used to exactly represent analytic solutions (i.e. (Q2)).
Previous sections show that concept (i) is indeed applicable to the radial equation, thus affirmatively answering (Q1). A radial scalar product was developed in Sec. IV. Methods for evaluation of the scalar product were developed in Sec. V. Connections between the scalar product and concept (ii) were discussed in Sec. VII. In particular, the scalar products connection with problem specific special functions (i.e. the confluent Heun polynomials) was detailed in Sec. VII.2. Here, we take a first step towards applying confluent Heun polynomials to the representation of analytic solutions to Teukolsky’s radial equation for QNMs. A non-classical feature of the confluent Heun polynomials means that, for now, only a preliminary and conceptual step is taken.
Recall that Teukolsky’s radial equation is a confluent Heun equation of the form,
$\mathcal{L}{{}_{\xi}}\;$ $=\;[\text{C}_{0}+\text{C}_{1}(1-\xi)]+\mathcal{D}_{\xi}\;.$ (101)
In Eq. (101), the transformed radial operator, $\mathcal{L}{{}_{\xi}}$, is simply rewritten such the differential part is explicitly identified as $\mathcal{D}_{\xi}$  (see Eq. 72). In Sec. VII.2, it was found that $\mathcal{D}_{\xi}$ generates confluent Heun polynomials, $y_{pk}(\xi)$, such that each $y_{pk}(\xi)$ satisfies a two parameter eigenvalue problem,
$\mathcal{D}{{}_{\xi}}\;|{{y}_{pk}}\rangle\;$ $=\;(\lambda_{pk}\;+\;\mu_{p}\,\xi)\;|{{y}_{pk}}\rangle\;,$ (102)
where, for a given polynomial order $p$, there are $p+1$ solutions, $y_{pk}$. This situation differs starkly from that of classical (e.g. Jacobi) polynomials, for which every single polynomial order maps to a single polynomial solution. In other words, for each polynomial order, classical polynomials have a simplex of solutions, while the confluent Heun polynomials have a multiplex.
The existence of a multiplex for each polynomial order prompts a nontrivial conceptual tension: The Jacobi polynomials’ recursion relationships play a key role in the angular problem (see Appx. B); however, there cannot exist similar recursion relationships between the confluent Heun polynomials, $y_{pk}$. This is because such recursion relationships require an orthogonal polynomial sequence, wherein each polynomial has a unique order (i.e. is a simplex) Chihara (2011). Further, since each confluent Heun polynomial multiplex is composed of polynomials of like order, the collection of all such multiplexes would appear to contain too many polynomials, meaning that the they allow for multiple equivalent representations (i.e. decompositions) of radial functions.
This does not, however, preclude the possibility that there are special combinations of confluent Heun polynomials allow for unique representations, and classical recursion relationships. While this possibility is perhaps too open ended to be fully pursued here, we may assume (somewhat adventurously) that such polynomials exist, and then contemplate implications for Teukolsky’s radial equation:
Suppose that solutions to the radial equation, $|{f}\rangle$, can be exactly represented in terms of orthonormal polynomials, $u_{n}$,
$|{f}\rangle\;=\;\sum_{p=0}^{\infty}a_{p}\,|{u_{p}}\rangle\;,$ (103a)
where it is additionally supposed that
$|{u_{p}}\rangle\;$ $=\;\sum_{k=0}^{p}\,\text{c}_{pk}\,|{{y}_{pk}}\rangle\;,$ (103b)
${\mathbb{I}}\;$ $=\;\sum_{p=0}^{\infty}\,|{u_{p}}\rangle\langle{u_{p}}|\;,$ (103c)
and
$\xi|{u_{p}}\rangle=\sigma_{p0}|{u_{p-1}}\rangle+\sigma_{p1}|{u_{p}}\rangle+\sigma_{p2}|{u_{p+1}}\rangle\;.$ (103d)
In Eqs. (103a-103d), despite not being classical, $u_{n}$ are constructed to have the essential properties of classical polynomials, namely completeness, orthonormality and three-term recursion Chihara (2011); Chen and Dai (2010). In this sense, we may think of $u_{n}$ as having canonical properties. This motivates their being referred to as canonical confluent Heun polynomials.
Further, in Eqs. (103a-103d), $a_{p}$, $c_{pk}$, $\sigma_{p0}$, $\sigma_{p1}$ and $\sigma_{p2}$ are assumed to be generally non-zero complex valued constants. Henceforth, $\sigma_{pj}$ will refer to complex valued constants that are assumed to be known if $u_{p}$ are known. In Eq. (103c), $\mathbb{I}$} is the identity operator for the space of all confluent Heun functions of the fractional radial coordinate, $\xi$. Equation (103c) requires that the expansion coefficients present in Eq. (103a) are $a_{p}={\langle{u_{p}}\,|\,{f}\rangle}$. The three-term-recursion relation given by Eq. (103d) is a universal feature of all (simplexical) orthogonal polynomial sequences, and it is a direct consequence of Eq. (103c) (i.e. completeness) along with the fact that ${\langle{u_{j}}\,|\,{\xi|u_{k}}\rangle}={\langle{\xi u_{j}}\,|\,{u_{k}}\rangle}={\langle{u_{j}}\,|\,{\xi u_{k}}\rangle}$ Chihara (2011).
Naively, if each order $p$ polynomial multiplex is a $p+1$ dimensional orthonormal basis, then it should be capable of exactly representing any order $p$ polynomial on the same domain Axler (2015). Thus it may be possible to construct $u_{p}$ such that Eqs. (103b-103d) hold.
In effect, Eq. (103) fully defines a solution ansatz. This solution ansatz may be used to investigate whether the scalar product and confluent Heun polynomials imply that Teukolsky’s radial equation is, like the angular equation, exactly tridiagonalizable. To this end, we may attempt to represent the radial equation in the basis defined by $\{u_{p}\}_{p=0}^{\infty}$.
In bra-ket notation, the transformed radial equation is
$\mathcal{L}{{}_{\xi}}\,|{f}\rangle\;$ $=\;A\,|{f}\rangle\;.$ (104)
Applying Eq. (103c) such that $\mathcal{L}_{\xi}$ is represented as a matrix, $\hat{L}$, and $|{f}\rangle$ a vector, $\vec{f}$, gives
$\sum_{p,p^{\prime}}\,|{u_{p^{\prime}}}\rangle$ $\langle{u_{p^{\prime}}}|\,\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle{\langle{u_{p}}\,|\,{f}\rangle}\;\;$ (105a)
$=\;A\,\sum_{p}\,|{u_{p}}\rangle{\langle{u_{p}}\,|\,{f}\rangle}\;,$
$\hat{L}\,\vec{f}\;$ $=\;A\,\vec{f}\;.$ (105b)
In Eq. (105b), the elements of $\hat{L}$ are $\langle{u_{p^{\prime}}}|\,\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$, and the elements of $\vec{f}$ are $\langle{u_{p}}\,|\,{f}\rangle$}. Henceforth, $p^{\prime}$ will label rows, and $p$ will label columns, meaning that $p\geq p^{\prime}$ denotes the upper triangle of $\hat{L}$, and $p\leq p^{\prime}$ its lower. To compute the matrix elements, $\langle{u_{p^{\prime}}}|\,\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$, the action of $\mathcal{L}{{}_{\xi}}$ on $|{u_{p}}\rangle$ must be considered in more detail.
Rewriting $\mathcal{L}{{}_{\xi}}$ in terms of $\mathcal{D}_{\xi}$ (see Eq. 101) allows $\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ to be expanded as
$\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle\;=\;(\text{C}_{0}+\text{C}_{1}[1-\xi])\,|{u_{p}}\rangle+\mathcal{D}_{\xi}\,|{u_{p}}\rangle\;.$ (106)
In Eq. (106), $\mathcal{D}_{\xi}\,|{u_{p}}\rangle$ can be rewritten using the definition of $u_{p}$ in terms of confluent Heun polynomials (Eq. 103b), and the two parameter eigenvalue relationship (Eq. 102). Doing so yields
$\mathcal{D}_{\xi}\,|{u_{p}}\rangle\;$ $=\;\sum_{k=0}^{p}\,\text{c}_{pk}\,\mathcal{D}_{\xi}\,|{{y}_{pk}}\rangle$ (107a)
$=\;\sum_{k=0}^{p}\,\text{c}_{pk}\,(\lambda_{pk}\,+\,\mu_{p}\,\xi)\,|{{y}_{pk}}\rangle\;,$ (107b)
$=\;|{v_{p}}\rangle\,+\,\mu_{p}\,\xi\,|{u_{p}}\rangle\;.$ (107c)
In going from Eq. (107b) to Eq. (107c), the term dependent on the confluent Heun eigenvalues, $\lambda_{pk}$, has been used to define a new polynomial,
$|{v_{p}}\rangle\;=\;\sum_{k=0}^{p}\,\text{c}_{pk}\,\lambda_{pk}\,|{{y}_{pk}}\rangle\;,$ (108)
and it has been noticed that the remaining term in Eq. (107b) is simply proportional to $|{u_{n}}\rangle$.
It follows from Eq. (107c) that $\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ may be rewritten as
$\mathcal{L}{{}_{\xi}}|{u_{p}}\rangle=(\text{C}_{0}+\text{C}_{1})|{u_{p}}\rangle+|{v_{p}}\rangle+(\mu_{p}-\text{C}_{1})\xi|{u_{p}}\rangle\;,$ (109)
and that the matrix element, $\langle{u_{p^{\prime}}}|\mathcal{L}{{}_{\xi}}|{u_{p}}\rangle$, is
$\langle{u_{p^{\prime}}}|\mathcal{L}{{}_{\xi}}|{u_{p}}\rangle\;=\;$ $(\text{C}_{0}+\text{C}_{1}){\langle{u_{p^{\prime}}}\,|\,{u_{p}}\rangle}$ (110a)
$+(\mu_{p}-\text{C}_{1})\langle{u_{p^{\prime}}}|\xi|{u_{p}}\rangle$ (110b)
$+{\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle}\;.$ (110c)
Since $|{u_{p}}\rangle$ are orthonormal (i.e. ${\langle{u_{p^{\prime}}}\,|\,{u_{p}}\rangle}=\delta_{p^{\prime}p}$), Eq. (110a) only contributes to diagonal matrix elements. Since $|{u_{p}}\rangle$ possess a three term recursion relation given by Eq. (103d), Eq. (110b) will only contribute when $|p^{\prime}-p|\leq 1$.
It happens that $\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle$} is not only non-zero when $|p^{\prime}-p|\leq 1$, but also restricted to the upper-triangle (i.e. $p\geq p^{\prime}$). This follows from the fact that $\mathcal{D}_{\xi}$ is self-adjoint w.r.t. the scalar product, meaning that
${\langle{u_{p^{\prime}}}\,|\,{\mathcal{D}_{\xi}\,u_{p}}\rangle}\;=\;{\langle{\mathcal{D}_{\xi}\,u_{p^{\prime}}}\,|\,{u_{p}}\rangle}\;.$ (111)
Using Eq. (107c) to expand both sides of Eq. (111) yields,
${\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle}-{\langle{v_{p^{\prime}}}\,|\,{u_{p}}\rangle}\;=\;(\mu_{p}-\mu_{p^{\prime}}){\langle{u_{p^{\prime}}}\,|\,{\xi|u_{p}}\rangle}\;.$ (112)
Equation (112) may be understood thusly: Since $u_{p^{\prime}}$ is a member of a complete orthonormal sequence, its scalar product with any order $p$ polynomial will be non-zero if $p\geq p^{\prime}$ and zero otherwise Chihara (2011). In other words, any order $p$ polynomial may be expressed as a linear combination of $u_{p^{\prime}}$, where $p^{\prime}\leq p$ is required since the order of interest is $p$. Thus $\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle$} and $\langle{v_{p^{\prime}}}\,|\,{u_{p}}\rangle$} only have non-zero values in the respective upper ($p\geq p^{\prime}$) and lower ($p\leq p^{\prime}$) triangle of $\hat{L}$. Further, $\xi\,u_{p}$ is a linear combination of $\{u_{p-1},u_{p},u_{p+1}\}$ due to Eq. (103d), so the matrix whose elements are $\langle{u_{p^{\prime}}}\,|\,{\xi|u_{p}}\rangle$} will be tridiagonal. Together, these points mean that the only way that Eq. (112) can hold, i.e. the only way that the upper and lower triangular structure of $\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle$} and $\langle{v_{p^{\prime}}}\,|\,{u_{p}}\rangle$} can be equal to the tridiagonal structure of $\langle{u_{p^{\prime}}}\,|\,{\xi|u_{p}}\rangle$}, is if $\{{\langle{u_{p^{\prime}}}\,|\,{v_{p}}\rangle}\}_{p^{\prime}p}$ is upper-triangular and tridiagonal. This means that $|{v_{p}}\rangle$ must be a linear combination of $u_{p-1}$ and $u_{p}$,
$|{v_{p}}\rangle=\sigma_{p3}|{u_{p-1}}\rangle+\sigma_{p4}|{u_{p}}\rangle\;,$ (113)
where $\sigma_{p3}={\langle{u_{p-1}}\,|\,{v_{p}}\rangle}$ and $\sigma_{p4}={\langle{u_{p}}\,|\,{v_{p}}\rangle}$ are complex valued constants which are determined by $c_{pk}$ (i.e. $u_{p}$) and the confluent Heun eigenvalues $\lambda_{pk}$.
Together, Eq. (103d), Eq. (110) and Eq. (113) allow $\mathcal{L}_{\xi}|{u_{p}}\rangle$ to be written as
$\mathcal{L}_{\xi}|{u_{p}}\rangle\;=\;\sigma_{p5}\,|{u_{p-1}}\rangle+\sigma_{p6}\,|{u_{p}}\rangle+\sigma_{p7}\,|{u_{p+1}}\rangle\;,$ (114)
where
$\sigma_{p5}\;$ $=\;(\mu_{p}-\text{C}_{1})\sigma_{p0}+\sigma_{p3}\;,$ (115a)
$\sigma_{p6}\;$ $=\;(\mu_{p}-\text{C}_{1})\sigma_{p1}+\sigma_{p4}+(\text{C}_{0}+\text{C}_{1})\;,$ (115b)
$\sigma_{p7}\;$ $=\;(\mu_{p}-\text{C}_{1})\sigma_{p2}\;.$ (115c)
The implication of Eqs. (114-115) is that, if the canonical polynomials $u_{n}$ can be determined, then Teukolsky’s radial operator is exactly tridiagonalizable.
IX Discussion
The primary results of this article are (i) the introduction of a radial scalar product for QNMs (Sec. IV), (ii) the development of QNM specific confluent Heun polynomials that may be defined for arbitrary frequency parameter (Sec. VII.2), and (iii) the demonstration that, in principle, the scalar product and proposed canonical confluent Heun polynomials enable the exact tridiagonalization of Teukolsky’s radial equation for QNMs (Sec. VIII).
It is also of note that the scalar product may be evaluated by direct integration with a complex valued radius, or by more standard analytic extension using confluent hypergeometric functions (See Sec. VII.1). While analytic extension was found to be the more practical route, the ability to construct a complex radial coordinate, $\xi(z)$, for evaluation of the scalar product also implies (via Eq. (30)) that the QNMs themselves are spatially bounded on $\xi(z)$. Since $\xi(z)$ must be constructed to coincide with Sturm-Liouville boundary conditions, a QNM on $\xi(z)$ would have physically relevant values at both the event horizon and spatial infinity.
Regarding the canonical polynomials, a clear next step is to investigate their computation. Like the Gram-Schmidt polynomials, which are unique once certain processing choices are made, one might expect that there are multiple constructions of the canonical polynomials. This a topic of ongoing research that is near conclusion London and Gurevich (2023a).
Similarly, the tridiagonalization of Teukolsky’s radial equation directly implies that its solutions (for fixed frequency parameter) are orthogonal. This would be directly analogous to the orthogonality of the spheroidal harmonics at fixed frequency (i.e. fixed “oblateness”) London (2023). It would also be compatible with the conclusions of complementary studies of linear perturbations of Kerr that have encountered a kind of QNM orthogonality Green et al. (2023).
An adjacent implication is that the radial functions are complete (i.e. isomorphic to canonical confluent Heun polynomials). If the radial functions are both orthogonal and complete for given value of the frequency parameter $\tilde{\omega}$, then this would strongly imply the existence of a biorthogonal dual for the radial solutions determined by the QNM frequencies. Again, this would be directly analogous to the QNM’s spheroidal harmonics, where each spheroidal harmonic requires input from a different QNM frequency London (2023). A consequence would be that the QNMs are “quasi-complete”, i.e. complete in the radial and angular domains, but not complete (specifically under- or over-complete) in the time domain. Investigation of these implications is also a topic of ongoing work that is near conclusion London and Gurevich (2023b).
Lastly, these ideas imply a potentially useful direction of research for e.g. BBH post-merger signals: rather than attempting to fit various QNMs’ within numerical data, it may be possible to directly project out individual QNMs’ time series using orthogonality. In the context of ongoing efforts in Numerical Relativity and signal modeling for future gravitational wave detectors, this may be a fruitful direction for future research.
X Acknowledgements
[…]
Appendix A Notation
This section lists explicit expressions for the main text’s schematic notation. The details here are provided for convenience and completeness.
In Boyer-Lindquist coordinates, Teukolsky’s master differential operator, $\mathcal{L}_{tr\theta\phi}$, is
$\mathcal{L}_{tr\theta\phi}\;\;=\;\left[\frac{(r^{2}+a^{2})^{2}}{\Delta(r)}-a^{2}\sin^{2}\theta\right]\frac{\partial^{2}}{\partial{t}^{2}}$ (116)
$+\frac{4Mar}{\Delta(r)}\frac{\partial^{2}}{\partial{t}\partial\phi}+\left[\frac{a^{2}}{\Delta(r)}-\frac{1}{\sin^{2}\theta}\right]\frac{\partial^{2}}{\partial\phi^{2}}$
$-\Delta(r)^{-s}\frac{\partial}{\partial{r}}\left(\Delta(r)^{s+1}\frac{\partial}{\partial{r}}\right)-\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta\frac{\partial}{\partial\theta}\right)$
$-2s\left[\frac{a(r-M)}{\Delta(r)}+\frac{i\cos\theta}{\sin^{2}\theta}\right]\frac{\partial}{\partial\phi}$
$-2s\left[\frac{M(r^{2}-a^{2})}{\Delta(r)}-r-ia\cos\theta\right]\frac{\partial}{\partial{t}}$
$+(s^{2}\cot^{2}\theta-s)\;,$
where
$\Delta(r)\;$ $=\;(r-r_{+})(r-r_{-})\;,$ (117a)
and
$r_{\pm}\;$ $=M\,\pm\,\sqrt{M^{2}-a^{2}}\;.$ (117b)
Upon separation via a separable ansatz or integral transform (see Sec. III), Teukolsky’s angular differential operator is encapsulated by the following expression,
$u\;=\;\cos(\theta)\;,$ (118a)
$\mathcal{L}_{u}=V_{S}(u)+\partial_{u}(1-u^{2})\partial_{u}\;,$ (118b)
$V_{S}(u)\;=\;s+ua\tilde{\omega}(ua\tilde{\omega}-2s)-\frac{(m+su)^{2}}{1-u^{2}}\;.$ (118c)
Similarly, Teukolsky’s radial differential operator is commonly written as
$\mathcal{L}{{}_{r}}\;$=\; $V_{R}(r)\;+\;\Delta(r)^{-s}\partial_{r}\left[\Delta(r)^{s+1}\partial_{r}\right]\;,$ (119)
where
$V_{R}(r)\;$ $=\;\frac{K(r)^{2}-2is(r-M)K(r)}{\Delta(r)}+\;4is\tilde{\omega}{}{r}$ (120a)
$\hskip 96.73918pt-a^{2}\tilde{\omega}{}^{2}+2am\tilde{\omega}{}$
$K(r)\;$ $=\;(r^{2}+a^{2})\,\tilde{\omega}{}-am\;.$ (120b)
The expressions above may also be found in Refs. Teukolsky and Press (1974); Leaver (1985); O’Sullivan and Hughes (2014).
Appendix B Solving Teukolsky’s angular equation with Jacobi polynomials
Teukolsky’s angular equation is
$\mathcal{L}_{u}\,S(u)\;=\;-\,AS(u)\;.$ (121)
In Eq. (121),
$u\;$ $=\;\cos(\theta)\;,$ (122a)
$\mathcal{L}_{u}$ $=V_{S}(u)+\partial_{u}(1-u^{2})\partial_{u}\;,$ (122b)
where the equation’s potential, $V_{S}$, is
$V_{S}(u)\;$ $=\;s+u\gamma(u\gamma-2s)-\frac{(m+su)^{2}}{1-u^{2}}\;$ (122c)
with
$\gamma\;$ $=\;a\tilde{\omega}\;$. (122d)
Like Teukolsky’s radial equation, Eq. (121) is a confluent Heun equation; however, Eq. (121) differs from the radial case in two ways of current relevance:
  • (i)
    Equation (121) has two regular singular points at $u\in\{-1,1\}$, and one irregular singular point at $u\rightarrow\infty$. The irregular singular point at $u\rightarrow\infty$ is not within the physical domain (i.e. there is no boundary condition defined there). The relevant asymptotic boundary conditions are essentially that $S(u)$ are regular when $u\in\{-1,1\}$.
  • (ii)
    The degeneration of Eq. (121) to a hypergeometric equation occurs at a physical limit ($a\rightarrow 0$), whereas the degeneration of the radial equation to a hypergeometric one occurs at a non-physical limit for QNMs ($\tilde{\omega}\rightarrow 0$).
Respectively, these differences mean that (i) global [^19] By global, one means a solution that applies to the entire domain. This is in contrast to local solutions, such as those use in the MST method, that are only valid in the vicinity of a singular point Mano et al. (1996); Ronveaux and Arscott (1995). solutions to the angular equation can be solved with high accuracy using non-confluent basis functions, and that (ii) the optimal basis functions may be constructed from physical solutions at the $a\rightarrow 0$ limit.
The construction of analytic solutions begins by finding a similarity transformation that results in the transformed potential being regular when $u\in[-1,1]$. The similarity transformation is equivalent to taking the solution ansatz,
$S(u)\;=\;\eta(u)\,g(u)\;,$ (123)
where,
$\eta(u)\;=\;e^{uk_{2}}\;(1-u)^{k_{0}}\;(1+u)^{k_{1}}\;,$ (124)
and then applying it to the angular equation thusly,
$\left[\eta(u)^{-1}\,\mathcal{L}{{}_{u}}\,\eta(u)\right]g(u)\;$ $=\;-\left[\eta(u)^{-1}A\,\eta(u)\right]g(u)\;$ (125a)
$\mathcal{L}{{}_{u}^{\prime}}\;$ $=\;\left[\eta(u)^{-1}\,\mathcal{L}{{}_{u}}\,\eta(u)\right]$ (125b)
$\mathcal{L}{{}_{u}^{\prime}}\,g(u)\;$ $=\;-\,A\,g(u)\;.$ (125c)
In Eq. (125a), the similarity transformation of $\mathcal{L}{{}_{u}}$ is explicitly performed. In Eq. (125b), the transformed operator, $\mathcal{L}{{}_{u}^{\prime}}$, is defined, and in Eq. (125c) its eigenvalue problem is stated.
The values of $k_{0}$, $k_{1}$ and $k_{2}$ that result in the potential of $\mathcal{L}{{}_{u}^{\prime}}$ being regular and linear on $u\in\{-1,1\}$ are,
$k_{0}\;$ $=\;\frac{1}{2}|m+s|\;,$ (126a)
$k_{1}\;$ $=\;\frac{1}{2}|m-s|\;,$ (126b)
$k_{2}\;$ $=\;\gamma\;.$ (126c)
Equivalently, $k_{0}$, $k_{1}$ and $k_{2}$ are the singular exponents that may be found via the method of Frobenius Courant and Hilbert (1989); Arfken et al. (2013); Pinchover and Rubinstein (2005). In effect, $k_{0}$ and $k_{1}$ remove divergences from the potential at the regular singular points, and $k_{2}$ removes the potential’s quadratic dependence on $u$. The appearance of the absolute values, e.g. $|m+s|$, ensures that $\eta$ is always finite when $u\in\{-1,1\}$, regardless of the values of $m$ and $s$. This is equivalent to enforcing asymptotic boundary conditions on $S(u)$. In Eq. (126c), $\gamma$ is also a viable solution for $k_{2}$; however, $k_{2}=+\gamma$ is chosen by convention Leaver (1985). The result for $\mathcal{L}{{}_{u}^{\prime}}$ has two schematic forms of relevance,
$\mathcal{L}_{u}^{\prime}\;=\;$ $(\text{a}_{0}+\text{a}_{1}\,u)$ (127a)
$+\;\left(\text{a}_{2}+\text{a}_{3}\,u+\text{a}_{4}\,[1-u^{2}]\right)\,\partial_{u}$ (127b)
$+\;\left(1-u^{2}\right)\partial^{2}_{u}\;,$ (127c)
and
$\mathcal{L}_{u}^{\prime}\;=\;$ $(\text{a}_{0}+\text{a}_{1}\,u)\,+\,\mathcal{D}^{\prime}_{u}\;.$ (128)
In Eq. (127), $\text{a}_{0}$ through $\text{a}_{4}$ are constant coefficients defined in Eqs. (129a-129e). They should not to be confused with the similarly labeled coefficients defined in the main text.
$\text{a}_{0}\;$ $=\;-k_{0}(2k_{1}+1)-k_{1}-\frac{1}{2}[m^{2}+s(s+2)]$ (129a)
$\;\;\;\;\;\,+\gamma^{2}+2\gamma(k_{1}-k_{0})$
$\text{a}_{1}\;$ $=\;-2\gamma(k_{0}+k_{1}+s+1)$ (129b)
$\text{a}_{2}\;$ $=\;-2(k_{0}-k_{1})$ (129c)
$\text{a}_{3}\;$ $=\;-2(k_{0}+k_{1}+1)$ (129d)
$\text{a}_{4}\;$ = $=\;2\gamma$ (129e)
Equation (128) defines $\mathcal{D}^{\prime}_{u}$ to be the differential part of Eq. (127) (i.e. Eqs. 127b and 127c). The operator, $\mathcal{D}^{\prime}_{u}$, is useful for understanding the eigenproblem (Eq. 125c) in the non-spinning limit (i.e. $\gamma=a\tilde{\omega}=0$), and for general spins.
In the non-spinning limit, $\mathcal{D}^{\prime}_{u}$ is simply $\mathcal{D}^{\text{Jac.}}_{u}$, the differential operator for which the Jacobi polynomials are eigenfunctions:
$\mathcal{D}^{\prime}_{u}|_{\gamma=0}\;$ $=\;\mathcal{D}^{\text{Jac.}}_{u}\;.$ (130)
The Jacobi operator is
$\mathcal{D}^{\text{Jac.}}_{u}\;$=\; $-(\,\alpha-\beta+(\alpha+\beta+2)\,u\,)\partial_{u}$ (131)
$+(1-u^{2})\partial^{2}_{u}\;,$
where
$\alpha\;$ = $=\;2\,k_{0}\,$, (132)
$\beta\;$ = $=\;2\,k_{1}\,.$ (133)
The simplest functions generated by $\mathcal{D}^{\text{Jac.}}_{u}$ are the Jacobi polynomials, $p_{n}(u)$. Their eigenvalue relationship is
$\mathcal{D}^{\text{Jac.}}_{u}\,p_{n}(u)\;$ $=\;-n(\alpha+\beta+n+1)\,p_{n}(u)\;.$ (134)
In Eq. (134), the eigenvalue is determined by the requirement that the series solution for $p_{n}(u)$ terminates after $n$ terms.
The zero-spin eigenvalue, $A|_{a=0}=A|_{\gamma=0}$, can be determined by requiring equivalence between the physical eigenvalue problem (Eq. 125c), and the Jacobi one (Eq. 134). This amounts to holding the non-derivative terms in each eigenvalue relation to be equivalent in the zero spin limit,
$(a_{0}+A)|_{\gamma\rightarrow 0}\;=\;n(\alpha+\beta+n+1)$ (135)
Equation (135) can be solved for the eigenvalue,
$A_{0}\;$ $=\;A|_{\gamma=0}$ (136a)
$=\;-\left(\ell-s\right)\left(\ell+s+1\right)\;.$ (136b)
In Eq. (136b), $A_{0}$ is the usual spin weighted spherical harmonic eigenvalue Newman and Penrose (1966); Leaver (1985); Fackerell and Crossman (2008), and index $\ell$ is related to the polynomial order, $n$, via
$\ell\;$ $=\;\ell_{\min}+n\,$, (137a)
$\ell_{\min}\;$ $=\;k_{0}+k_{1}\;.$ (137b)
For non-zero spin, term by term comparison of $\mathcal{D}^{\prime}_{u}$ and $\mathcal{D}^{\text{Jac.}}_{u}$ yields
$\mathcal{D}^{\prime}_{u}\;$ $=\;\mathcal{D}^{\text{Jac.}}_{u}+\text{a}_{4}(1-u^{2})\partial_{u}\;.$ (138)
Equation (138), in effect, casts $\mathcal{D}^{\prime}_{u}$ as a deformation of the Jacobi operator. This implies that it is natural to consider the spheroidal functions $g(u)$ to be deformations of the Jacobi polynomials. In particular, we will take $g(u)$ to be exactly represented by Jacobi polynomials,
$g(u)\;=\;\sum_{n=0}^{\infty}\,a_{n}\,p_{n}(u)\;.$ (139)
Equation (139) is justified not only because the Jacobi polynomials are known to be complete, but also because $g(u)$ and $p_{n}(u)$ satisfy the same asymptotic boundary conditions on the same domain.
Equation (139) is an analytic solution for $g(u)$ that is motivated by the structure of Teukolsky’s angular equation. In effect, Eq. (139) is a spectral decomposition of $g(u)$ into Jacobi moments, $a_{n}$. Thus, finding the values of each $a_{n}$ is equivalent to finding $g(u)$.
To do so, the Jacobi polynomials’ recursive and orthogonality properties must be used. The relevant recursion rules for Jacobi polynomials are
$(1-u^{2})\,\partial_{u}p_{n}\;$ $=\;\text{b}_{0}\,p_{n-1}+\text{b}_{1}\,p_{n}+\text{b}_{2}\,p_{n+1}\;,$ (140a)
$u\,p_{n}\;$ $=\;\text{b}_{3}\,p_{n-1}+\text{b}_{4}\,p_{n}+\text{b}_{5}\,p_{n+1}\;.$ (140b)
Normalization of the Jacobi polynomials is defined with respect to the weighted inner product,
${\langle{p_{n}}\,|\,{p_{n^{\prime}}}\rangle}\;$ $=\;\int_{-1}^{1}p_{n}(u)\,p_{n^{\prime}}(u)\;\text{W}^{\text{Jac.}}(u)\,\text{d}u$ (141a)
$=\;\text{b}_{6}\,\delta_{n^{\prime}n}\;,$ (141b)
where the Jacobi weight function, $\text{W}^{\text{Jac.}}(u)$, is
$\text{W}^{\text{Jac.}}\;=\;(1-u)^{\alpha}(1+u)^{\beta}\;.$ (142)
In Eqs. (140) and (141), $\text{b}_{0}$ through $\text{b}_{6}$ are $n$ dependent constants defined in Eqs. (143a-143g).
$\text{b}_{0}(n)\;$ $=\;-\frac{2(\alpha+n)(\beta+n)(\alpha+\beta+n+1)}{(\alpha+\beta+2n)(\alpha+\beta+2n+1)}$ (143a)
$\text{b}_{1}(n)\;$ $=\;-\frac{2n(\beta-\alpha)(\alpha+\beta+n+1)}{(\alpha+\beta+2n)(\alpha+\beta+2n+2)}$ (143b)
$\text{b}_{2}(n)\;$ $=\;\frac{2n(n+1)(\alpha+\beta+n+1)}{(\alpha+\beta+2n+1)(\alpha+\beta+2n+2)}$ (143c)
$\text{b}_{3}(n)\;$ $=\;-\frac{2(\alpha+n)(\beta+n)}{(\alpha+\beta+2n)(\alpha+\beta+2n+1)}$ (143d)
$\text{b}_{4}(n)\;$ $=\;-\frac{\alpha^{2}-\beta^{2}}{(\alpha+\beta+2n)(\alpha+\beta+2n+2)}$ (143e)
$\text{b}_{5}(n)\;$ $=\;-\frac{2(n+1)(\alpha+\beta+n+1)}{(\alpha+\beta+2n+1)(\alpha+\beta+2n+2)}$ (143f)
$\text{b}_{6}(n)\;$ $=\;\frac{\Gamma(n+\alpha+1)\Gamma(n+\beta+1)}{(\alpha+\beta+2n+1)\Gamma(n+1)\Gamma(n+\alpha+\beta+1)}$ (143g)
Observe that $\text{W}^{\text{Jac.}}(u)$ is very similar to the scaling function $\eta(u)$ introduced in Eq. (124). This can be understood as resulting from the fact that the transformation present in Eq. (125a) is conformal, meaning that
$\int_{-1}^{1}p_{n}(u)\,p_{n^{\prime}}$ $(u)\,\text{W}^{\text{Jac.}}(u)\,\text{d}u$ (144a)
$\int_{-1}^{1}([p_{n}(u)\eta(u)$ $]\,[p_{n^{\prime}}(u)\eta(u)])|_{\gamma=0}\,\text{d}u$ (144b)
There is no weight function in Eq. (144b) because Eq. (122b) is explicitly in Sturm-Liouville form (with weight function 11Arfken et al. (2013). The integrands of Eqs. (144a) and (144b) must be equivalent, meaning that $\text{W}^{\text{Jac.}}(u)$ and $\eta(u)|_{\gamma=0}$ are related,
$\text{W}^{\text{Jac.}}(u)\;=\;\eta(u)^{2}|_{\gamma=0}\;.$ (145)
In general, it can be shown that if $f_{1}$ satisfies the second order linear differential equation $\mathcal{L}_{1}f_{1}=0$, where $\mathcal{L}$ is self-adjoint with weight function $\lambda_{1}$, then $f_{2}=\eta^{-1}f_{1}$ will satisfy $\mathcal{L}_{2}g=0$, where $\mathcal{L}_{2}=\eta^{-1}\mathcal{L}_{1}\eta$ is self-adjoint with new weight function $\lambda_{2}=\eta^{2}\lambda_{1}$.
Equations (140-143) can be used to determine the values of $a_{n}$ as follows. Given the scalar product, Eq. (141a), we will use bra-ket notation to simplify to representation of various steps. In that notation, the completeness of the Jacobi polynomials is expressed as
$\sum_{n=0}^{\infty}\,|{p_{n}}\rangle\langle{p_{n}}|\;=\;{\mathbb{I}}\;.$ (146)
In Eq. (146), $\mathbb{I}$} is the identity operator for all functions that have the same asymptotic behavior as the Jacobi polynomials on $u\in[-1,1]$, and the polynomials $|{p_{n}}\rangle$ are implicitly normalized so that ${\langle{p_{n}}\,|\,{p_{n}}\rangle}=1$. The bra-ket notation form of the physical eigenvalue relation is,
$\mathcal{L}{{}_{u}^{\prime}}\,|{g}\rangle\;$ $=\;-\,A\,|{g}\rangle\;.$ (147)
From this perspective, the Jacobi moments, $a_{n}$, are simply the components of $g(u)$ in the basis of Jacobi polynomials. This is made explicit by using Eq. (146) to express both sides of Eq. (147) in the Jacobi polynomial basis,
$\sum_{n,n^{\prime}}|{p_{n^{\prime}}}\rangle\langle{p_{n^{\prime}}}|\mathcal{L}{{}_{u}^{\prime}}|{p_{n}}\rangle{\langle{p_{n}}\,|\,{g}\rangle}\;$ $=\;-\,A\,\sum_{n}|{p_{n}}\rangle{\langle{p_{n}}\,|\,{g}\rangle}\;.$ (148)
Note that the right hand side of Eq. (148) is equivalent to that of Eq. (138) (i.e. $a_{n}={\langle{p_{n}}\,|\,{g}\rangle}$). Equation (148) may be recast as a simple matrix equation
$\hat{L}\,\vec{g}\;$ $=\;-A\,\vec{g}\;,$ (149)
where $\hat{L}$ and $\vec{g}$ are defined from Eq. (148) as
$\hat{L}\;$ $=\;\{\langle{p_{n^{\prime}}}|\mathcal{L}{{}_{u}^{\prime}}|{p_{n}}\rangle\}_{n,n^{\prime}}\;,$ (150a)
$\vec{g}\;$ $=\;\{{\langle{p_{n}}\,|\,{g}\rangle}\}_{n}\;.$ (150b)
Thus, taking a finite dimensional truncation of $\hat{L}$ allows $\vec{g}$ (i.e. $a_{n}$) to be determined using standard linear algebraic methods Axler (2015); Virtanen et al. (2020).
The essential feature of this approach is that, starting with Eq. (149), the recursion and orthogonality relationships, Eqs. (140-143), can be used to show that $\hat{L}$ is at most tridiagonal. In particular, when $\mathcal{L}{{}_{u}^{\prime}}$ acts on $|{p_{n}}\rangle$, only lower to upper diagonal terms contribute,
$\mathcal{L}{{}_{u}^{\prime}}|{p_{n}}\rangle\;=\;\sigma_{n}\,|{p_{n-1}}\rangle+\mu_{n}\,|{p_{n}}\rangle+\nu_{n}\,|{p_{n+1}}\rangle\;.$ (151)
In Eq. (151), the $n$ dependent constants, $\sigma_{n}$, $\mu_{n}$, and $\nu_{n}$ are
$\sigma_{n}\;$ $=\;\frac{\text{b}_{6}(n-1)}{\text{b}_{6}(n)}[\text{a}_{1}\text{b}_{3}(n)+\text{a}_{4}\text{b}_{0}(n)]$ (152a)
$\mu_{n}\;$ $=\;\text{a}_{0}+\text{a}_{1}\text{b}_{4}(n)+\text{a}_{4}\text{b}_{1}(n)$ (152b)
$\quad\quad\quad\quad\quad\quad-n(\alpha+\beta+n+1)$
$\nu_{n}\;$ $=\;\frac{\text{b}_{6}(n+1)}{\text{b}_{6}(n)}[\text{a}_{1}\text{b}_{5}(n)+\text{a}_{4}\text{b}_{2}(n)]$ (152c)
For any $N$ dimensional truncation of $\hat{L}$, say $\hat{L}_{N}$, Eq. (152) provides the matrix elements,
$\hat{L}_{N}\;$ $=\;\left(\begin{array}[]{ccccc}\mu_{0}&\nu_{0}&0&\cdots&0\\ \sigma_{1}&\mu_{1}&\nu_{1}&\ddots&0\\ 0&\sigma_{2}&\mu_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\nu_{N-1}\\ 0&0&\cdots&\sigma_{N}&\mu_{N}\end{array}\right)\;.$ (158)
The use of Eq. (158) to determine angular functions $S(u)=\eta(u)g(u)$, and their eigenvalues $A(u)$, amounts to determining the eigenvectors and eigenvalues of $\hat{L}_{N}$. Since this method explicitly estimates the point spectra of $\hat{L}$, it is a purely spectral method (e.g. not a generalized spectral method) for the construction of analytic solutions to Teukolsky’s angular equation.
This construction facilitates the following conclusions: Since the transformed angular equation degenerates to the Jacobi equation when $\gamma=0$, the spheroidal functions $g(u)$ may be exactly identified with Jacobi polynomials in that limit. For general values of $\gamma$, the fact that $\hat{L}$ is tridiagonal means that for every $g(u)$ associated with index $\ell$, there always exists a unique Jacobi polynomial with that same index London (2023). In this way, the QNM index $\ell$ is inherently linked to a Jacobi polynomial of order $n=\ell-\ell_{\text{min}}$.
Appendix C The confluent Heun equation in natural general form
The confluent Heun equation has many forms that differ with respect to transformations of the independent and dependent variables Ronveaux and Arscott (1995). Here, we briefly outline the connection between Teukolsky’s radial equation and the confluent Heun equation in natural general form.
Perhaps the most referenced (e.g. Refs. Zarrinkamar and Hortaçsu (2018); Leaver (1986); Cook (2020)) form of confluent Heun equation is,
$\mathcal{L}^{(1)}_{\text{cH}}\,y(z)\;=\;0\;,$ (159a)
where
$\mathcal{L}^{(1)}_{\text{cH}}\;=\;(\alpha$ $\nu z-\sigma)$ (159b)
$\;+\,$ $\left[\gamma(z-1)+z(\delta+\nu(z-1))\right]\,\partial_{z}$
$\;+\,$ $(z-1)z\,\partial^{2}_{z}\;.$
Equation (159) may be derived from the Heun equation (which has $4$ regular singular points) by confluence of two singular points (see e.g. Arfken et al. (2013)) into a single irregular one. Equation (159a) is simply a definition of the confluent Heun equation in terms of its associated linear differential operator, $\mathcal{L}^{(1)}_{z}$. In Eq. (159b), $\alpha$, $\nu$, $\sigma$, $\gamma$ and $\delta$ are complex valued parameters, and $z$ is nominally real valued.
Equation (159), is the confluent Heun equation in what’s called nonsymmetrical canonical form Ronveaux and Arscott (1995); Cook and Zalutskiy (2014); Fiziev (2009b). It happens that this is actually not the most general form of the confluent Heun equation. As a result, a naive comparison between Eq. (159) and Teukolsky’s radial equation may generate unnecessary confusion. To facilitate the upcoming discussion, we will henceforth use $r$ rather than $z$, and we will discuss forms of the confluent Heun equation by referring to the related differential operators, $\mathcal{L}^{(i)}_{\text{cH}}$, such that
$\mathcal{L}^{(i)}_{\text{cH}}\,y\;=\;0\;.$ (160)
The most general form of the confluent Heun differential operator with an irregular singular point at infinity is (see e.g. Ref. Ronveaux and Arscott (1995))
$\mathcal{L}^{(2)}_{\text{cH}}\;$=\; $\left({\alpha_{0}}+{\alpha_{1}}r+{\alpha_{2}}r^{2}+\frac{{\alpha_{4}}}{r-{r_{-}}}+\frac{{\alpha_{3}}}{r-{r_{+}}}\right)$ (161)
$+\;\left({\alpha_{5}}+{\alpha_{6}}r+{\alpha_{7}}r^{2}\right)\,\partial_{r}$
$+\;(r-{r_{-}})(r-{r_{+}})\,\partial^{2}_{r}\;.$
In Eq. (161), $\mathcal{L}^{(2)}_{\text{cH}}$ corresponds to the confluent Heun equation in what is called natural general form Ronveaux and Arscott (1995).
In Eq. (161), $\mathcal{L}^{(2)}_{\text{cH}}$ has been formatted to ease comparison with Teukolsky’s radial differential operator. In particular, the two regular singular points have been placed at $r_{+}$ and $r_{-}$, just as they are in the case of Teukolsky’s radial differential operator (Eq. 119), which can be formatted as
$\mathcal{L}{{}_{r}}\;$=\; $\left({\text{A}_{0}}+{\text{A}_{1}}r+\text{A}_{2}\,r^{2}+\frac{{\text{A}_{3}}}{r-{r_{-}}}+\frac{{\text{A}_{4}}}{r-{r_{+}}}\right)$ (162)
$+\;({\text{A}_{5}}+{\text{A}_{6}}r)\,\partial_{r}$
$+\;(r-{r_{-}})(r-{r_{+}})\,\partial_{r}^{2}\;.$
For completeness, parameters $\text{A}_{1}$ through $\text{A}_{6}$ are defined in terms of $_{+}, rr_{-}$ and $s$ as
$\text{A}_{0}\;=$ $\;\tilde{\omega}{}({r_{-}}+{r_{+}})(\tilde{\omega}{}({r_{-}}+{r_{+}})-is)$ (163a)
$\text{A}_{1}\;=$ $\;\tilde{\omega}{}(\tilde{\omega}{}({r_{-}}+{r_{+}})+2is)$ (163b)
$\text{A}_{2}\;=$ $\;\tilde{\omega}{}^{2}$ (163c)
$\text{A}_{3}\;=$ $\;\frac{{r_{-}}^{2}\left(m^{2}+4{r_{-}}^{2}\tilde{\omega}{}^{2}\right)}{{r_{-}}-{r_{+}}}$ (163d)
$-\;{r_{-}}\left(m^{2}+{r_{-}}\tilde{\omega}{}(3{r_{-}}\tilde{\omega}{}+is)\right)$
$+\;\frac{4m{r_{-}}^{2}\tilde{\omega}{}\sqrt{{r_{-}}{r_{+}}}}{{r_{+}}-{r_{-}}}+ims\sqrt{{r_{-}}{r_{+}}}$
+ $+\;2m{r_{-}}\tilde{\omega}{}\sqrt{{r_{-}}{r_{+}}}-{r_{-}}{r_{+}}\tilde{\omega}{}({r_{-}}\tilde{\omega}{}+is)$
$\text{A}_{4}\;=$ $\;-\frac{{r_{+}}^{2}\left(m^{2}+4{r_{+}}^{2}\tilde{\omega}{}^{2}\right)}{{r_{-}}-{r_{+}}}$ (163e)
$-\;{r_{+}}\left(m^{2}+{r_{+}}\tilde{\omega}{}(3{r_{+}}\tilde{\omega}{}+is)\right)$
$+\;\frac{4m{r_{+}}^{2}\tilde{\omega}{}\sqrt{{r_{-}}{r_{+}}}}{{r_{-}}-{r_{+}}}+ims\sqrt{{r_{-}}{r_{+}}}$
+ $+\;2m{r_{+}}\tilde{\omega}{}\sqrt{{r_{-}}{r_{+}}}-{r_{-}}{r_{+}}\tilde{\omega}{}({r_{+}}\tilde{\omega}{}+is)$
$\text{A}_{5}\;=$ $\;(s+1)(-({r_{-}}+{r_{+}}))$ (163f)
$\text{A}_{6}\;=$  $\;2(s+1)\;.$ (163g)
The fact that Teukolsky’s radial operator is a special case of Heun’s confluent operator is clear by comparing Eq. (161) to Eq. (162). The confluent Heun nature of the operator is not changed under transformations of the dependent or independent variables, but such transformations can impact the physical nature of solutions. For example, application of the QNM boundary conditions (see e.g. Eqs. 11-13) has the effect of transforming Eq. (162) into a different version of Eq. (161). However, if the solution of the transformed equation (e.g. Eq. (15)) is to be consistent with QNM boundary conditions, then it cannot have asymptotic behavior that affects a transformation into yet another version of the confluent Heun equation. Application of a compactified coordinate (see Eq. 10) has the effect of changing the coordinate position of singular points, but not their regular or irregular nature.